### v1

# Heegaard Floer Surgery Formula and Cosmetic Surgeries

Abstract.
Two Dehn surgeries on a knot are called cosmetic if they yield homeomorphic three-manifolds. We show for a certain family of null-homologous knots in any closed orientable three-manifold, if the knot admits cosmetic surgeries with a pair of positive surgery coefficients, then the coefficients are both greater than $1$. In addition, for this family of knots, we show that $1/q$ Dehn surgery for $q$ at least $2$ is not homeomorphic to the original three-manifold. The proofs of these results use the mapping cone formula for the Heegaard Floer homology of Dehn surgery in terms of the knot Floer homology of the knot; we provide a new proof of this formula for integer surgeries in $\operatorname{Spin}^{c}$ structures with nontorsion first Chern class.
1. Introduction
Heegaard Floer homology is a powerful invariant of closed oriented three-manifolds introduced by Ozsváth and Szabó [OS04b]. It has been used to prove many results about three-manifold topology, especially Dehn surgery. In fact, the Heegaard Floer homology of any surgery on a knot can be computed from the knot Floer complex of the knot [OS08, OS10, Zem23].
In this paper, we will use Heegaard Floer homology to study cosmetic surgeries on a certain class of knots in arbitrary closed oriented three-manifolds. We recall the definition of cosmetic surgeries.
Definition 1.1.
If two Dehn surgeries on a knot yield homeomorphic manifolds, then these two surgeries are cosmetic. If in addition, the two surgeries are orientation-preserving homeomorphic, then these two surgeries are purely cosmetic.
Cosmetic surgeries are very rare; in fact, the following cosmetic surgery conjecture has been proposed:
Conjecture 1.2 ([Kir95] Problem 1.81).
Suppose $K$ is a knot in a closed oriented three-manifold $Y$ such that $Y-K$ is irreducible and not homeomorphic to the solid torus. If two different Dehn surgeries on $K$ are purely cosmetic, then there is a homeomorphism of $Y-K$ which takes one slope to the other.
The existence of cosmetic surgeries has been ruled out in many cases for knots in $S^{3}$. For instance, Ozsváth and Szabó proved that if $K\subset S^{3}$ admits cosmetic rational surgeries, then either the resulting manifold is an $L$-space or the surgery slopes have opposite signs [OS10]. Much more is known when considering purely cosmetic surgeries. Hanselman proved that any pair of purely cosmetic Dehn surgeries on knots in $S^{3}$ must have slopes that are ±$\pm 2$ or ±$\pm 1/q$ for some integer $q$ [Han19].
Less is known for knots in arbitrary three-manifolds. Ni found an obstruction to cosmetic surgeries on knots in three-manifolds satisfying a certain condition on their Thurston norms [Ni10].
The main theorem in this paper is an analogue of Ozsváth and Szabó’s result. We consider a class of knots which were shown by Ni to have special properties on their knot Floer homology.
Definition 1.3.
Let $Y$ be a closed, oriented three-manifold, and $K\subset Y$ a null-homologous knot. We say that $K$ is blow-up unknotted if there exists a taut Seifert surface $F$ for $K$, some natural number $N$, and a properly-embedded smooth disk $G$ inside $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ with $\partial G=K\times\{0\}$ and $[G]=\iota_{*}[F]\in H_{2}(X,K\times\{0\})$ for $\iota:(Y,K)\to(X,K\times\{0\})$ the inclusion.
Remark.
The property of $K$ being blow-up unknotted implies that $K$ is null homotopic.
For example, by Kirby calculus, $K$ is a blow-up unknotted knot if $K$ is obtained from the unknot in $Y$ by a sequence of crossing changes that change a positive crossing to a negative one.
Theorem 1.4.
Let $K\subset Y$ be a blow-up unknotted knot, and suppose $0<r<s$ are distinct positive rational numbers with the property that $Y_{r}(K)\cong\pm Y_{s}(K)$. Then $r,s>1$ or $K$ is the unknot.
We also provide progress on the knot complement problem for blow-up unknotted knots.
Theorem 1.5.
Let $K\subset Y$ be a nontrivial blow-up unknotted knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $g(K)\geq 2$, then $Y_{1}(K)\ncong\pm Y$.
The proofs of these theorems rely on a mapping cone formula that relates the Heegaard Floer homology of Dehn surgery on null-homologous knots in arbitrary closed, oriented three-manifolds with the knot Floer chain complex. This result was proven for Heegaard Floer homology in $\operatorname{Spin}^{c}$ structures whose first Chern classes are torsion by Ozsváth and Szabó, and the formula in nontorsion $\operatorname{Spin}^{c}$ structures for zero surgeries was proven by Ni [OS08, OS10, Ni20]. A recent result by Zemke provides a very general surgery formula via an equivalence of certain Lagrangians with local systems in the Fukaya category [Zem23]. Perhaps of independent interest, we provide a conceptually-simpler proof of the integer surgery formula for nontorsion $\operatorname{Spin}^{c}$ structures, which generalizes Ozsváth and Szabó’s and Ni’s results.
Acknowledgements The author would like to thank Yi Ni for advising him in this project.
2. Preliminaries on Heegaard Floer Homology
In this section, we establish some notation and collect some results that will be needed in later sections.
Let $W:Y_{1}\to Y_{2}$ be an oriented cobordism, with $Y_{1},Y_{2}$ connected. Given $\mathfrak{s}\in\operatorname{Spin}^{c}(W)$, there is a homomorphism
$F_{W,\mathfrak{s}}^{\circ}:HF^{\circ}(Y_{1},\mathfrak{s}|Y_{1})\to HF^{\circ}(Y_{2},\mathfrak{s}|Y_{2}),$
induced by a chain map
$f_{W,\mathfrak{s}}^{\circ}:CF^{\circ}(Y_{1},\mathfrak{s}|Y_{1})\to CF^{\circ}(Y_{2},\mathfrak{s}|Y_{2}).$
Among other properties, these maps satisfy a conjugation invariance. For $\mathfrak{t}$ a $\operatorname{Spin}^{c}$ structure (in either three or four dimensions), let $\overline{\mathfrak{t}}$ denote its conjugate. There is an isomorphism $\mathfrak{I}_{Y}:HF^{\circ}(Y,\mathfrak{t})\to HF^{\circ}(Y,\overline{\mathfrak{t}})$.
Theorem 2.1 ([OS06] Theorem 3.6).
Let $W$ be a cobordism from $Y_{1}$ to $Y_{2}$ as above. Then
$F_{W,\mathfrak{s}}^{\circ}=\mathfrak{I}_{Y_{2}}\circ F_{W,\overline{\mathfrak{s}}}^{\circ}\circ\mathfrak{I}_{Y_{1}}.$
Given a chain map $f:A\to B$, let $MC(f)$ be the mapping cone of $f$, namely, $MC(f)=A\oplus B$ with the differential
$\partial=\begin{pmatrix}\partial_{A}&0\\ f&\partial_{B}\end{pmatrix}.$
Theorem 2.2 ([OS08] Theorem 3.1).
Let $Y$ be a closed, oriented three-manifold, $K\subset Y$ a null-homologous knot. Fix an integer $n$ and a positive integer $m$. Then, there is a long exact sequence
$\cdots\to HF^{+}(Y_{n}(K))\to HF^{+}(Y_{m+n}(K))\to\bigoplus^{m}HF^{+}(Y)\to\cdots$
and also a corresponding exact sequence using $\widehat{HF}$ in place of $HF^{+}$. There are ${\mathbb{Z}}[U]$-equivariant chain maps
$f_{1}^{+}:CF^{+}(Y_{n}(K))\to CF^{+}(Y_{m+n}(K))$
$f_{2}^{+}:CF^{+}(Y_{m+n}(K))\to\bigoplus^{m}CF^{+}(Y)$
$f_{3}^{+}:\bigoplus^{m}CF^{+}(Y)\to CF^{+}(Y_{n}(K))$
inducing the maps in the long exact sequence, and ${\mathbb{Z}}[U]$-equivariant quasi isomorphisms
$\phi^{+}:CF^{+}(Y_{n}(K))\to MC(f_{2}^{+})$
$\psi^{+}:MC(f_{2}^{+})\to CF^{+}(Y_{n}(K)).$
Knot Floer homology associates to a knot $K$ in a closed, oriented three-manifold $Y$ a ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered ${\mathbb{Z}}[U]$-complex $C=CFK^{\infty}(Y,K)$. Write the two factors of the ${\mathbb{Z}}\oplus{\mathbb{Z}}$ filtration as $(i,j)$. Let $S$ be a region in the plane such that for all $(i,j)\in S$ and all $(i^{\prime},j^{\prime})\geq(i,j)$, we have that $(i^{\prime},j^{\prime})\in S$. Define $A_{s}^{+}=C\{\max(i,j-s)\geq 0\}$ and $B^{+}=C\{i\geq 0\}$ as in Ozsváth-Szabó. There are two canonical chain maps $v_{s}^{+}:A_{s}^{+}\to B^{+}$ and $h_{s}^{+}:A_{s}^{+}\to B^{+}$, where $v_{s}^{+}$ is the vertical projection, and $h_{s}^{+}$ first projects $A_{s}^{+}$ to $C\{j\geq s\}$, then maps to $C\{j\geq 0\}$ via $U^{s}$, and finally maps to $B^{+}$ via a chain homotopy equivalence $C\{j\geq 0\}\to C\{i\geq 0\}$ [OS08].
Let $K\subset Y$ be null-homologous. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, and let $n\in{\mathbb{Z}}$. The space of $\operatorname{Spin}^{c}$ structures over $Y_{n}(K)$ which are $\operatorname{Spin}^{c}$-cobordant to $\mathfrak{t}$ over the cobordism $W_{n}^{\prime}(K)$ is identified with ${\mathbb{Z}}/n{\mathbb{Z}}$ by $\mathfrak{u}\mapsto i$, where $\mathfrak{s}$ is an extension of $\mathfrak{u}$ over $W_{n}^{\prime}(K)$ with the property that
$\langle c_{1}(\mathfrak{s}),[\hat{F}]\rangle-n\equiv 2i\mod 2n.$
For $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ denote the corresponding summand of
$CF^{+}(Y_{n}(K))$. Let $A_{s,\mathfrak{t}}^{+},B_{\mathfrak{t}}^{+}$ denote the corresponding chain complexes associated to the knot filtration on $CF^{+}(Y,\mathfrak{t})$ [OS08].
Let
${\mathbb{A}}_{i,\mathfrak{t}}^{+}=\prod_{s\in{\mathbb{Z}}|s\equiv i\mod n}A_{s,\mathfrak{t}}^{+}\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{\mathfrak{t}}^{+}=\prod_{s\in{\mathbb{Z}}|s\equiv i\mod n}B_{\mathfrak{t}}^{+}$
and let $D_{n,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{i,\mathfrak{t}}^{+}$ be the map
$D_{n,i,\mathfrak{t}}^{+}(\{a_{s}\})=\{b_{s}\},$
where here
$b_{s}=h_{s-n}^{+}(a_{s-n})+v_{s}^{+}(a_{s}).$
Let ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ denote the mapping cone of $D_{n,i,\mathfrak{t}}^{+}$ [OS08].
Taking the direct sum of the above chain complexes over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/n{\mathbb{Z}}$ gives ${\mathbb{A}}^{+}$, ${\mathbb{B}}^{+}$, and the map $D_{n}^{+}:{\mathbb{A}}^{+}\to{\mathbb{B}}^{+}$. Let ${\mathbb{X}}^{+}(n)$ denote the mapping cone of $D_{n}^{+}$.
Let $K\subset Y$ be a null-homologous knot and let $F$ be a Seifert surface for $K$ in the homology class $\phi$. Given $n>0$ and $t\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ be the natural two-handle cobordism, and let $\widehat{F}_{n}\subset W_{n}^{\prime}(K)$ be the closed surface obtained by capping of $\partial F$ with a disk. Let $\widehat{\phi}_{n}=[\widehat{F}_{n}]\in H_{2}(W_{n}^{\prime}(K))$. Let $k$ be an integer satisfying $k\equiv t\mod n$ and $|k|\leq n/2$. Let $\mathfrak{x}_{k},\mathfrak{y}_{k}\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ satisfy $\mathfrak{t}_{k}|Y=\mathfrak{n}_{k}|Y=\mathfrak{s}$ and
$\langle c_{1}(\mathfrak{x}_{k}),\hat{\phi}_{n}\rangle=2k-n\hskip 14.22636pt\langle c_{1}(\mathfrak{y}_{k}),\hat{\phi}_{n}\rangle=2k+n,$
and let $\mathfrak{s}_{t}=\mathfrak{t}_{k}|Y_{n}(K)=\mathfrak{n}_{k}|Y_{n}(K)$.
Theorem 2.3 ([OS08] Theorem 2.3).
When $n\geq 2g(F)$, the chain complex
$CF^{+}(Y_{n}(K),\mathfrak{s}_{t})$ is represented by the chain complex $A_{k}^{+}$, in the sense that there is an isomorphism
$\Psi_{n}^{+}:CF^{+}(Y_{n}(K),\mathfrak{s}_{t})\to A_{k}^{+}.$
Moreover, the following squares commute:
{tikzcd}
{tikzcd}
In the case of rational surgeries, let $p,q$ be a pair of relatively prime integers. For $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/p{\mathbb{Z}}$, denote
${\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}=\prod_{r\in{\mathbb{Z}}}(r,A_{\lfloor(i+pr)/q\rfloor,\mathfrak{t}}^{+}(K))\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}=\prod_{r\in{\mathbb{Z}}}(r,B_{\mathfrak{t}}^{+}).$
Let $D_{p/q,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}$ be the map
$D_{p/q,i,\mathfrak{t}}^{+}\{(r,a_{r})\}=\{(r,b_{r})\},$
where
$b_{r}=v_{\lfloor(i+pr)/q\rfloor}^{+}(a_{r})+h_{\lfloor(i+p(r-1))/q\rfloor}^{+}(a_{r-1}).$
Let ${\mathbb{X}}_{p/q,i,\mathfrak{t}}^{+}$ denote the mapping cone of $D_{p/q,i,\mathfrak{t}}^{+}$ [OS10].
As before, direct summing over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/p{\mathbb{Z}}$ gives ${\mathbb{A}}_{p/q}^{+}$, ${\mathbb{B}}_{p/q}^{+}$, and $D_{p/q}^{+}:{\mathbb{A}}_{p/q}^{+}\to{\mathbb{B}}_{p/q}^{+}$. Let ${\mathbb{X}}_{p/q}^{+}$ denote the mapping cone of $D_{p/q}^{+}$. In the case where $p/q$ is clear from the context, we will drop the subscript $p/q$ from the notation.
3. Integer Surgeries
In this section, we provide a proof of the mapping cone formula for integer surgeries in the case where the $\operatorname{Spin}^{c}$ structure has nontorsion first Chern class. The formula for torsion $\operatorname{Spin}^{c}$ structures was originally proven by Ozsváth and Szabó [OS10]. Our proof combines ingredients from their argument as well as Ni’s proof of the zero surgery formula.
Let $K\subset Y$ be a null-homologous knot. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$ with nontorsion first Chern class.
Fix $n>0$. The long exact sequence in Theorem 2.2 decomposes over $\operatorname{Spin}^{c}$ structures on $Y$. Setting $m=nk$ for large $k$, it also decomposes over $\operatorname{Spin}^{c}$ structures on $Y_{n}(K)$, i.e. given $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ where
(3.1) $f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k+1)}(K))$ $|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k+1)\equiv 2s\mod 2n(k+1),$
$\text{and }s\equiv i\mod n\},$
$CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ |\ s\equiv i\mod n}CF^{+}(Y_{n(k+1)}(K),s,\mathfrak{t}).$
A strengthening of the proof of Proposition 3.5 in [Ni20] shows the following:
Proposition 3.2 ([Ni20] Proposition 3.5).
Let $p=\max\{k+g(F),-2k+g(F)\}$. Suppose that $n\geq\max\{p,2g(F)\}$, $\mathfrak{s}\in\mathcal{X}_{k}\setminus\{\mathfrak{x}_{k},\mathfrak{y}_{k}\},$ then the map $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through the map
$U^{n-p}:\operatorname{HF}^{-}(Y_{n}(K),k,\mathfrak{t})\circlearrowright.$
Moreover, if $c_{1}(\mathfrak{t})$ is nontorsion, then $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{+}$ factorizes through
$U^{n-p}:\operatorname{HF}^{+}(Y_{n}(K),k,\mathfrak{t})\circlearrowright.$
Proof.
The proof is exactly the same as in [Ni20], except we note that
$\frac{l(l+1)n}{2}-lk-g\geq n-p$
whenever $l\neq 0,-1$. From here, it follows that $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through $U^{n-p}$. ∎
Theorem 3.3.
For each $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, the mapping cone ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ of
$D_{n,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{i,\mathfrak{t}}^{+}$
is identified with $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
Proof when $n>0$.
Recall that $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ as in Equation 3.1. Since $c_{1}(\mathfrak{t})$ is nontorsion, $H_{*}(A_{i}^{+})$ is a finitely generated abelian group. So $U^{p}|H_{*}(A_{i}^{+})=0$ when $p$ is greater than a constant $C_{1}$ independent of $n,k$. By Theorem 2.3, this implies that $U^{p}|HF^{+}(Y_{n(k+1)}(K),i,\mathfrak{t})=0$ when $p>C_{1}$, $n(k+1)>2g$.
Recall that we choose representatives
$-\frac{n(k+1)}{2}\leq s<\frac{n(k+1)}{2}$
for classes in ${\mathbb{Z}}/n(k+1){\mathbb{Z}}$. Suppose $s\geq 0$. Then for $l=\max\{s+g,-2s+g\}$, we have
$l\leq\frac{n(k+1)}{2}+g\leq n(k+1)$
for large $k$. Proposition 3.2 implies that $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}=0$ when $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$.
Now for $s<0$, let $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then by conjugation invariance, there is a chain equivalence
$F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}\cong F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{+}.$
If $\mathfrak{s}_{s}=\mathfrak{x}_{s}+r\operatorname{PD}[\widehat{F}]$, then
$\overline{\mathfrak{s}_{s}}=\overline{\mathfrak{x}}_{-s}-(r+1)\operatorname{PD}[\widehat{F}],$
in particular, $\overline{\mathfrak{s}_{s}}$ is not one of the two special $\operatorname{Spin}^{c}$ structures and $-s>0$. By the above argument, $F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{+}=0$, hence also $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}=0$.
The conclusion follows by Theorem 2.3. ∎
3.1. Proof of Theorem for negative surgery coefficients
Continue to assume $n>0$, only we consider $Y_{-n}(K)$. In this case, let $W^{\prime}_{n(k-1)}(K)$ be the cobordism from $Y_{n(k-1)}(K)$ to $Y$. For each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/n{\mathbb{Z}}$,
$HF^{+}(Y_{-n}(K),i,\mathfrak{t})\cong H_{*}\left(MC\left(f_{2,i,\mathfrak{t}}^{+}:CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to\bigoplus^{nk}CF^{+}(Y,\mathfrak{t})\right)\right),$
where
$f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k-1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k-1)}(K))$ $|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k-1)\equiv 2s\mod 2n(k-1),$
$\text{and }s\equiv i\mod n\},$
$CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k-1){\mathbb{Z}}\ |\ s\equiv i\mod n}CF^{+}(Y_{n(k-1)}(K),s,\mathfrak{t}).$
Consider the summand $CF^{+}(Y_{n(k-1)}(K),s,\mathfrak{t})$ for
$-\frac{n(k-1)}{2}\leq s<\frac{n(k-1)}{2},\hskip 14.22636pts\equiv i\mod n.$
Then, as above, if $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then $F_{W^{\prime}_{n(k-1)}(K),\mathfrak{s}_{s}}^{+}=0$. The conclusion again follows by Theorem 2.3.
4. Rational Surgeries
In this section, we sketch the proof of the rational surgery formula:
Theorem 4.1.
Let $K\subset Y$ be a null-homologous knot, and let $p,q$ be a pair of relatively prime integers. Then, for each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/p{\mathbb{Z}}$, there is an isomorphism of ${\mathbb{Z}}[U]$-modules
$H_{*}(\mathbb{X}_{i,\mathfrak{t},p/q}^{+})\cong HF^{+}(Y_{p/q}(K),i,\mathfrak{t}).$
Ozsváth and Szabó stated and proved this theorem in the case where $\mathfrak{t}$ has torsion first Chern class [OS10]. In this case, they showed this isomorphism also preserves ${\mathbb{Z}}$-grading. Their argument still holds when the $\operatorname{Spin}^{c}$ structure is nontorsion, as long as grading information is ignored.
By Kirby calculus, $Y_{p/q}(K)$ can be realized by surgery with coefficient $a$ inside the knot $K\#O_{q/r}\subset Y\#(-L(q,r))$, where $a$ is the greatest integer smaller than or equal to $p/q$, $a=\lfloor p/q\rfloor$, and $p/q=a+r/q$, and $O_{q/r}\subset-L(q,r)$ is the knot which is obtained by viewing one component of the the Hopf link as a knot inside the lens space $-L(q,r)$ thought of as $-q/r$ surgery on the other component of the Hopf link [OS10]. Thus, one obtains the rational surgery formula by proving a formula for Morse surgeries on rationally null-homologous knots and then applying a Kunneth principle for connected sums.
Analogously to the case of integer surgeries on null-homologous knots, for $K\subset Y$ a rationally null-homologous knot, $\lambda$ a framing on $K$, and $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, Ozsváth and Szabó define the chain map $D_{\mathfrak{s}}^{+}:{\mathbb{A}}_{\mathfrak{s}}^{+}(Y,\overline{K})\to{\mathbb{B}}_{\mathfrak{s}}^{+}(Y,\overline{K})$. Let ${\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$ be the mapping cone of $D_{\mathfrak{s}}^{+}$ [OS10].
Theorem 4.2 (Morse surgery formula [OS10]).
Let $K\subset Y$ be a rationally null-homologous knot, and let $\lambda$ be a framing on $K$. Let $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, then
$HF^{+}(Y_{\lambda}(K),\mathfrak{s})\cong H_{*}({\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$
as ${\mathbb{Z}}[U]$-modules.
We recall the definition of $U$-knots from [OS10].
Definition 4.3.
A $U$-knot is a knot in a rational homology three-sphere $Y$ with the property that the induced filtration of $CF^{\infty}(Y,\mathfrak{s})$ is trivial. More precisely, given any $\xi\in\underline{\operatorname{Spin}^{c}}(Y,K)$, $CFK^{\infty}(Y,K,\xi)$ is chain homotopy equivalent via a relatively ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered homotopy to the chain complex $R$ which is a free, rank one ${\mathbb{Z}}[U,U^{-1}]$ module with the trivial differential.
Proposition 4.4.
If $K_{2}\subset Y_{2}$ is a $U$-knot, then for each $\xi_{1}\in\underline{\operatorname{Spin}}^{c}(Y_{1},\overline{K}_{1})$ and $\mathfrak{s}_{2}\in\operatorname{Spin}^{c}(Y_{2})$, there is some $\xi_{2}\in\underline{\operatorname{Spin}^{c}}(Y_{2})$ representing $\mathfrak{s}_{2}$, with the property that
$CFK^{\infty}(Y_{1},\overline{K_{1}},\xi_{1})\cong CFK^{\infty}(Y_{1}\#Y_{2},\overline{K_{1}}\#\overline{K_{2}},\xi_{1}\#\xi_{2})$
as ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered chain complexes [OS10].
Lemma 4.5.
$O_{q/r}\subset-L(q,r)$ is a $U$-knot. There is an affine identification $\phi$ of the $\operatorname{Spin}^{c}$ structures on $\widehat{HFK}(L(q,r),O_{q,r})$ with ${\mathbb{Z}}$ such that
$\widehat{HFK}(L(q,r),O_{q,r},\phi(i))\cong\begin{cases}{\mathbb{Z}}&\text{if $0\leq i\leq q-1$,}\\ 0&\text{otherwise.}\end{cases}$
[OS10]
Let $K\subset Y$ be a null-homologous knot. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, let $H^{2}(Y,K,\mathfrak{t})\cong{\mathbb{Z}}$ denote the subgroup of $H^{2}(Y,K)$ corresponding to $\mathfrak{t}$, and also $H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})\cong{\mathbb{Z}}$.
Lemma 4.6.
Let $K\subset Y$ be a null-homologous knot. Under the connected sum
$(Y,K)\#(-L(q,r),O_{q,r})\to(Y\#-L(q,r),K\#O_{q,r}),$
the following diagram commutes:
{tikzcd}
where $f(x,y)=qx+y$. Moreover, under this correspondence, if $K_{\lambda}$ is the pushoff of $K$ with respect to the integral framing $a$, then $\operatorname{PD}[K_{\lambda}]$ represents the element $p\in{\mathbb{Z}}\cong H^{2}(Y\#-L(q,r),K\#O_{q,r})$, where $p/q=a+r/q$.
Proof.
Follows as in [OS10]. ∎
With these ingredients, the proof of Theorem 4.1 follows as in [OS10].
5. Cosmetic Surgeries
In this section, we prove our main results. For simplicity, we consider coefficients in the field ${\mathbb{F}}={\mathbb{F}}_{2}$. Throughout, we let $K\subset Y$ be a null-homologous knot in a closed, oriented three-manifold. We will obstruct the result of Dehn surgeries on $K$ from being homeomorphic to each other using the rank of $\widehat{HF}(Y_{r}(K))$ for various positive rational numbers $r$.
We state some lemmas that will be useful in calculating the homology of the mapping cones for Dehn surgery.
The following lemma was proven in [Ni20].
Lemma 5.1.
If $s\leq t$, then $*\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{t})_{*}$ and $*\operatorname{Im}(\widehat{h}_{s})_{*}\supset\operatorname{Im}(\widehat{h}_{t})_{*}$.
Lemma 5.2.
There is an equivalence between $\widehat{v}_{s}$ and $\widehat{h}_{-s}$ for all $s$.
This equivalence is induced by changing the roles of the basepoints $w$ and $z$ in the doubly-pointed Heegaard diagram.
Lemma 5.3.
Let $g(K)$ be the genus of $K$, then for all $s\geq g(K)$, $\widehat{v}_{s}$, $\widehat{h}_{-s}$ induce isomorphisms on homology. For all $s\geq g(K)+1$, $\widehat{v}_{-s}$ and $\widehat{h}_{s}$ are nullhomotopic.
Lemma 5.4 ([Ni20] Lemma 4.1).
Let $K\subset Y$ be a null-homologous knot. Suppose that $F$ is a taut Seifert surface for $K$. Let $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ for some $N$, and let $\iota:(Y,K)\to(X,K\times\{0\})$ be the inclusion map. Let $G\subset X$ be a properly-embedded smooth connected surface with $\partial G=K\times\{0\}$, $[G]=\iota_{*}[F]\in H^{2}(X,K\times\{0\})$, and $g(G)<g(F)$. Then
$\operatorname{Im}(v_{k}^{+})_{*}\supset\operatorname{Im}(h_{k}^{+})_{*},$
when $k\geq g(G)$. Moreover, if $G$ is a disk, then
$(v_{0}^{+})_{*}=(h_{0}^{+})_{*}$
Again, call a knot where such $(X,G)$ for $G$ a disk exists a blow-up unknotted knot. We also consider the unknot itself to be a blow-up unknotted knot.
The following result and its proof are analogous to [OS10] Proposition 9.7.
Proposition 5.5.
Let $K\subset Y$ be a null-homologous knot. Then
$g(K)=\max(0,\{s\in{\mathbb{Z}}\ |\ (\widehat{v}_{s-1})_{*}\text{ is not an isomorphism}\}).$
In particular, $K$ is the unknot if and only if $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$.
The proof makes use of the $CF^{\delta}$ variant of Heegaard Floer homology defined in [OS08]. For any nonnegative integer $\delta$, Denote by $CF^{\delta}(Y)$ the subcomplex of $CF^{+}(Y)$ which is the kernel of multiplication by $U^{\delta+1}$. Similarly, let $A_{s}^{\delta}$, $B^{\delta}$ denote the subcomplexes of $A_{s}^{+}$, $B^{+}$, respectively, which are the kernel of multiplication by $U^{\delta+1}$.
Proof.
Let $g=g(K)$, by Lemma 5.3,
$g(K)\geq\max(0,\{s\in{\mathbb{Z}}\ |\ (\widehat{v}_{s-1})_{*}\text{ is not an isomorphism}\}).$
If $g(K)=0$, i.e. $K$ is the unknot, then $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Otherwise, if $g(K)>0$, it suffices to show $(\widehat{v}_{g-1})_{*}$ is not an isomorphism. Consider the short exact sequence
(5.6) $0\to C\{i<0,j\geq g-1\}\to C\{i\geq 0\text{ or }j\geq g-1\}\xrightarrow{v_{g-1}^{+}}C\{i\geq 0\}\to 0$
The middle term is $A_{g-1}^{+}$, and $C\{i\geq 0\}=B^{+}$. Also,
$H_{*}(C\{i<0,j\geq g-1\})\cong H_{*}(C\{(-1,g-1)\}\cong H_{*}(C\{0,g\})=\widehat{HFK}(K,g)\neq 0.$
Therefore, $(v_{g-1}^{+})_{*}$ is not an isomorphism.
Suppose for the sake of contradiction that $(\widehat{v}_{g-1})_{*}$ is an isomorphism. Suppose also that $(v_{g-1}^{\delta})_{*}$ is an isomorphism for some $\delta\geq 0$. By [OS04a] Theorem 4.4, the diagram in Figure 1 commutes and the rows are exact, so by the five lemma, $(v_{g-1}^{\delta+1})_{*}$ is also an isomorphism. By induction, $(v_{g-1}^{\delta})_{*}$ is an isomorphism for all $\delta$. Thus, by Lemma 2.7 in [OS08], $(v_{g-1}^{+})_{*}$ is an isomorphism, which is a contradiction.
{tikzcd}
Figure 1.
Lemma 5.7.
Let $K\subset Y$ be a blow-up unknotted knot, $p,q>0$ relatively prime, and let $x\in H_{*}(\widehat{A}_{j/q})$ for some $j$. If $\lfloor j/q\rfloor\geq 0$, there exists $\alpha\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ such that $(\widehat{D}_{p/q})_{*}(x+\alpha)=(\widehat{v}_{j/q})_{*}(x)$ (where we think of $H_{*}(\widehat{B}_{j/q})$ as including into $H_{*}(\widehat{\mathbb{B}})$). Similarly, if $\lfloor j/q\rfloor\leq 0$, there exists $\alpha\in H_{*}(\oplus_{i\leq j-p}\widehat{A}_{i/q})$ such that $(\widehat{D}_{p/q})_{*}(x+\alpha)=(\widehat{h}_{j/q})_{*}(x)$.
Proof.
For $s\geq 0$, we have $*\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s+1})_{*}\subset\dots$, and for $s\leq 0$, $*\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s-1})_{*}\subset\dots$. Thus, in the first case, there exists $y_{1}\in H_{*}(\widehat{A}_{(j+p)/q})$ such that $(\widehat{h}_{j/q})_{*}(x)=(\widehat{v}_{(j+p)/q})_{*}(y_{1})$. If $(\widehat{h}_{(j+p)/q})_{*}(y_{1})=0$, then stop and set $\alpha=y_{1}$, otherwise keep going. Inductively, given $y_{n}$ in $H_{*}(\widehat{A}_{(j+np)/q})$, there exists $y_{n+1}\in H_{*}(\widehat{A}_{(j+(n+1)p)/q})$ such that $(\widehat{h}_{(j+np)/q})_{*}(y_{n})=(\widehat{v}_{(j+(n+1)p)/q})_{*}$. By Lemma 5.3, this process eventually terminates and gives an element $\alpha=\sum_{n\geq 1}y_{n}\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ with the desired condition.
Similarly, in the case where $\lfloor j/q\rfloor\leq 0$, since $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\subset\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, we can inductively add cancellation terms to the left. ∎
Using the mapping cone formula and the above lemmas, in the case where $K$ is blow-up unknotted, we may now calculate the rank of $\widehat{HF}(Y_{r}(K))$ for any positive rational number $r$.
Lemma 5.8.
Let $K\subset Y$ be a blow-up unknotted knot, let $p,q>0$ coprime. Let $b=\operatorname{rank}(\widehat{B})=\operatorname{rank}\widehat{HF}(Y)$. Then
(5.9) $\begin{split}\operatorname{rank}\widehat{HF}(Y_{p/q}(K))&=q(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &+2q\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})+2t_{K}^{p/q}-pb,\end{split}$
where
(5.10) $t_{K}^{p/q}$ $=\sum_{j=0}^{p-1}\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right).$
Note that if $Y=S^{3}$, Equation 5.9 agrees with Equation 40 in [OS10]. Indeed, define $\nu(K)$ as in [OS10]. If $\nu(K)>0$, then
$\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right)=\begin{cases}1&j\geq q\nu(K)\text{ and }j<p-q(\nu(K)-1)\\ 0&\text{otherwise.}\end{cases}$
Thus,
$t_{K}^{p/q}=\max(0,p-(2\nu(K)-1)q).$
Also, for all $s\geq 0$,
$\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\begin{cases}\operatorname{rank}H_{*}(\widehat{A}_{s})_{*}&s<\nu(K)\\ \operatorname{rank}H_{*}(\widehat{A}_{s})_{*}-1&s\geq\nu(K),\end{cases}$
so the two equations agree.
If $\nu(K)=0$, then $t_{K}^{p/q}=p$ and $\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})_{*}-1$ for all $s\geq 0$, so again the two equations agree.
Proof.
By Theorem 4.1, $\widehat{HF}(Y_{p/q}(K))\cong H_{*}(\widehat{\mathbb{X}}_{p/q})$, the homology of the mapping cone of $\widehat{D}_{p/q}$ from $\widehat{{\mathbb{A}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{A}}_{i}$ to $\widehat{{\mathbb{B}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{B}}_{i}$. We abuse notation to let $\widehat{A}_{j/q}$ denote $\widehat{A}_{\lfloor j/q\rfloor}$, and also $\widehat{B}_{j/q}$, $\widehat{v}_{j/q}$, $\widehat{h}_{j/q}$ similarly. So
$\widehat{\mathbb{A}}=\bigoplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\prod_{r\in{\mathbb{Z}}}\widehat{A}_{(i+pr)/q}=\prod_{j\in{\mathbb{Z}}}\widehat{A}_{j/q}$
has $q$ copies of $\widehat{A}_{s}$ for each $s\in{\mathbb{Z}}$, and under the map $\widehat{v}$, each $\widehat{A}_{j/q}$ corresponds to a copy of $\widehat{B}$ in $\widehat{\mathbb{B}}$, which we shall denote $\widehat{B}_{j/q}$. Then $\widehat{h}_{j/q}$ maps $\widehat{A}_{j/q}$ to $\widehat{B}_{(j+p)/q}$. Thus, for each $s\in{\mathbb{Z}}$, there are $p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s-1}$ and $q-p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s}$.
The finiteness of $H_{*}(\widehat{\mathbb{X}}_{p/q})$ is guaranteed by Lemma 5.3; in fact, it implies that $\widehat{\mathbb{X}}_{p/q}$ is isomorphic to a truncated version of the mapping cone
$\widehat{\mathbb{X}}_{p/q}^{c}=MC(\widehat{D}_{p/q}^{c}:\widehat{\mathbb{A}}_{p/q}^{c}\to\widehat{\mathbb{B}}_{p/q}^{c}),$
for some large $c$, where
$\widehat{\mathbb{A}}_{p/q}^{c}=\bigoplus_{j=-qc+1}^{qc-1}\widehat{A}_{j/q}=\bigoplus_{s=-c+1}^{c-1}\bigoplus_{i=0}^{q-1}\widehat{A}_{s},\hskip 14.22636pt\widehat{\mathbb{B}}_{p/q}^{c}=\bigoplus_{j=-qc+p+1}^{qc-1}\widehat{B}_{j/q},$
and $\widehat{D}_{p/q}^{c}$ is the restriction of $\widehat{D}_{p/q}$ to $\widehat{A}_{p/q}^{c}$, with the exception that for any $x\in\widehat{A}_{j/q}$ with $-qc+1\leq j\leq-qc+p$, $\widehat{D}_{p/q}^{c}(x)=\widehat{h}_{j/q}(x)$, and for any $x\in\widehat{A}_{j/q}$ with $qc-p\leq j\leq qc-1$, $\widehat{D}_{p/q}^{c}(x)=\widehat{v}_{j/q}(x)$. Note that it suffices to take $c\geq g+p/q+1$, and
$\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})=q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s}),\hskip 14.22636pt\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})=qb(2c-1)-pb.$
We construct a basis for $*\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}$. For $j\geq 0$, let
$*\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$
be a basis of $*\ker(\widehat{v}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose an $\alpha_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\alpha_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$. For $j<0$, let
$*\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{h}_{j/q})_{*}}$
be a basis of $*\ker(\widehat{h}_{j/q})_{*}$. For each such $x_{j/q}^{i}$, choose a $\beta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\beta_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$. Lastly, for $10\leq j\leq p-$ 1, let
$\{y_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right)}$
be a basis of the subspace of $H_{*}(\widehat{A}_{j/q})$ that is mapped isomorphically to $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$ under $(\widehat{v}_{j/q})_{*}$. Choose $z_{j/q}^{i}\in H_{*}(\widehat{A}_{(j-p)/q})$ such that $(\widehat{v}_{j/q})_{*}(y_{j/q}^{i})=(\widehat{h}_{(j-p)/q})_{*}(z_{j/q}^{i})$. Choose $\gamma_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ and $\delta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{y}_{j/q}^{i}=y_{j/q}^{i}+z_{j/q}^{i}+\gamma_{j/q}^{i}+\delta_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$.
The set
$S=\{\widetilde{x}_{j/q}^{i}\}\cup\{\widetilde{y}_{j/q}^{i}\}$
is linearly independent, and it is finite by Lemma 5.3. Moreover, it spans $*\ker(\widehat{D}_{p/q})_{*}$: let $*z\in\ker(\widehat{D}_{p/q})_{*}$ be nonzero, and suppose first that $z\in H_{*}(\oplus_{k\geq 0}\widehat{A}_{k/q})$. Let $z_{j/q}$ denote the term of $z$ in $H_{*}(\widehat{A}_{j/q})$, and let
$j=\min\{k\geq 0\ |\ z_{k/q}\neq 0\}.$
Then $z\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, so $*z_{j/q}\in\operatorname{Span}\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$: otherwise if $*z_{j/q}\notin\ker(\widehat{v}_{j/q})_{*}$, then $*z\notin\ker(\widehat{D}_{p/q})_{*}$ since $z\in H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$. Thus,
$z-\sum_{i\in J}\widetilde{x}_{j/q}^{i}\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k>j}\widehat{A}_{k/q})$
[…]
$z=\sum_{k\in I}\sum_{i_{k}\in J_{k}}\widetilde{x}_{k/q}^{i_{k}}$
[…]
$z=\sum_{k\in I}\sum_{i_{k}\in J_{k}}\widetilde{x}_{k/q}^{i_{k}}$
[…]
Now for the general case $*z\in\ker(\widehat{D}_{p/q})_{*}$, we claim that for each $10\leq j\leq p-$ 1,
$z_{j/q}\in\operatorname{Span}\left(\{y_{j/q}^{i}\}\cup\{x_{j/q}^{i}\}\right).$
Indeed, suppose this is not true for some $10\leq j\leq p-$ 1, then $*(\widehat{v}_{j/q})_{*}(z_{j/q})\notin\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, so $(\widehat{D}_{p/q})_{*}(z)\neq 0$, a contradiction. Thus,
$z-\sum_{j=0}^{p-1}\left(\sum_{i\in J_{j}}\widetilde{y}_{j/q}^{i}+\sum_{i\in J_{j}^{\prime}}\widetilde{x}_{j/q}^{i}\right)$ $\in\left(\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq p}\widehat{A}_{k/q})\right)$
$\hskip 14.22636pt+\left(\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k<0}\widehat{A}_{k/q})\right).$
By the above argument, $z$ is in the span of $S$.
Therefore, $S$ is a basis for $*\ker(\widehat{D}_{p/q})_{*}$, so since $*\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}\ker(\widehat{h}_{-s})_{*}$,
$\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}=q\operatorname{rank}\ker(\widehat{v}_{0})_{*}+2q\sum_{s=1}^{g-1}\operatorname{rank}\ker(\widehat{v}_{s})_{*}+t_{K}^{p/q}.$
Since $*\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})-\operatorname{rank}\ker(\widehat{v}_{s})_{*}$, $\operatorname{rank}H_{*}(\widehat{A}_{s})=b$ for all $|s|\geq g$, and $*\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}=\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})-\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})+\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*}$ for some large $c$, it follows that
$*\displaystyle\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$ $=b(2c-1)q-pb-q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s})$
$\hskip 14.22636pt+q\operatorname{rank}\ker(\widehat{v}_{0})_{*}+2q\sum_{s=1}^{g-1}\operatorname{rank}\ker(\widehat{v}_{s})_{*}+t_{K}^{p/q}$
$=q(b-\operatorname{rank}H_{*}(\widehat{A}_{0})+\operatorname{rank}\ker(\widehat{v}_{0})_{*})$
$\hskip 14.22636pt+2q\sum_{s=1}^{g-1}(b-\operatorname{rank}H_{*}(\widehat{A}_{s})+\operatorname{rank}\ker(\widehat{v}_{s})_{*})+t_{K}^{p/q}-pb$
$=q(b-\operatorname{rank}(\widehat{v}_{0})_{*})+2q\sum_{s=1}^{g-1}(b-\operatorname{rank}(\widehat{v}_{s})_{*})+t_{K}^{p/q}-pb,$
and therefore Equation 5.9 holds.
Now we are ready to prove our main theorems. The main ingredient for each result is Equation 5.9.
Theorem 5.11.
Let $K\subset Y$ be a blow-up unknotted knot, and suppose $0<r<s$ are distinct positive rational numbers with the property that $Y_{r}(K)\cong\pm Y_{s}(K)$. Then $r,s>1$ or $K$ is the unknot.
Proof.
Since $H_{1}(Y_{\pm p/q}(K))\cong H_{1}(Y)\oplus{\mathbb{Z}}/p{\mathbb{Z}},$ we can fix $p>0$ throughout. First, let $q$ be relatively prime to $p$ and $q\geq p$. Then $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}=\operatorname{Im}(\widehat{v}_{0})_{*}$ for all $j=0,\dots,p-1$. Thus $*t_{K}^{p/q}=p\operatorname{rank}(\widehat{v}_{0})_{*}$. Therefore, according to equation 5.9, for positive integral $q$ with $0<p/q\leq 1$, $\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ is a nondecreasing function of $q$. The function is strictly increasing unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Hence, by Proposition 5.5, $K$ is the unknot.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result is true whenever $0<r<s\leq 1$.
Now suppose $q,q^{\prime}$ are relatively prime to $p$ and $0<q<p<q^{\prime}$. Then
$t_{K}^{p/q}-t_{K}^{p/q^{\prime}}$ $*\displaystyle\leq\sum_{j=0}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-p\operatorname{rank}(\widehat{v}_{0})_{*}$
$*\displaystyle=\sum_{j=q}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-(p-q)\operatorname{rank}(\widehat{v}_{0})_{*}$
$\leq(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*}).$
Therefore,
$\operatorname{rank}\widehat{HF}(Y_{p/q^{\prime}}(K))$ $-\operatorname{rank}\widehat{HF}(Y_{p/q}(K))=(q^{\prime}-q)(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})$
$+(q^{\prime}-q)\left(2\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})\right)$
$+2t_{K}^{p/q^{\prime}}-2t_{K}^{p/q}$
$\geq 2(q^{\prime}-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})-2(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})$
$\geq 0,$
with equality only if $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$, hence $K$ is the unknot by Lemma 5.5.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result holds. ∎
In the case where $p$ is $1$ or $2$, the following corollary is immediate, since there do not exist distinct positive integers $q,q^{\prime}$ coprime to $p$ with $p/q$ and $p/q^{\prime}$ both greater than $1$.
Corollary 5.12.
Let $K\subset Y$ be a blow-up unknotted knot, and suppose $0<q^{\prime}<q$ are distinct positive numbers with the property that $Y_{p/q}(K)\cong\pm Y_{p/q^{\prime}}(K)$ for $p\in\{1,2\}$, then $K$ is the unknot.
Theorem 5.13.
Let $K\subset Y$ be a nontrivial blow-up unknotted knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $g(K)\geq 2$ or $\operatorname{rank}H_{*}(\widehat{A}_{0})>b$, then $Y_{1}(K)\ncong\pm Y$.
Proof.
Let $q\geq 2$, then by Equation 5.9,
$\operatorname{rank}\widehat{HF}(Y_{1/q}(K))\geq(b-\operatorname{rank}(\widehat{v}_{0})_{*})(q-1)+\operatorname{rank}(\widehat{v}_{0})_{*}+q\operatorname{rank}\ker(\widehat{v}_{0})_{*}\geq b,$
and it is strictly greater unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$. Hence, by Proposition 5.5, $K$ is the unknot.
If $q=1$, then
$\operatorname{rank}\widehat{HF}(Y_{1}(K))\geq\operatorname{rank}(\widehat{v}_{0})_{*}+\operatorname{rank}\ker(\widehat{v}_{0})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{0})\geq b,$
and it is strictly greater unless $\operatorname{rank}H_{*}(\widehat{A}_{0})=b$ and $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 1$, which implies $g(K)\leq 1$.
Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result holds. ∎

### v2

@@
 Heegaard Floer Surgery Formula and Cosmetic Surgeries
-Abstract.
+Abstract
 Two Dehn surgeries on a knot are called cosmetic if they yield homeomorphic three-manifolds.
 We show for a certain family of null-homologous knots in any closed orientable three-manifold, if the knot admits cosmetic surgeries with a pair of positive surgery coefficients, then the coefficients are both greater than $1$.
-In addition, for this family of knots, we show that $1/q$ Dehn surgery for $q$ at least $2$ is not homeomorphic to the original three-manifold.
+In addition, for this family of knots, we show that +1/q+1/q Dehn surgery is not homeomorphic to the original three-manifold.
 The proofs of these results use the mapping cone formula for the Heegaard Floer homology of Dehn surgery in terms of the knot Floer homology of the knot; we provide a new proof of this formula for integer surgeries in $\operatorname{Spin}^{c}$ structures with nontorsion first Chern class.
-1. Introduction
+Keywords: Heegaard Floer Homology, knot Floer homology, Dehn surgery, cosmetic surgery
+1 Introduction
 Heegaard Floer homology is a powerful invariant of closed oriented three-manifolds introduced by Ozsváth and Szabó [OS04b].
@@
 Conjecture 1.2 ([KIR95] Problem 1.81).
-Suppose $K$ is a knot in a closed oriented three-manifold $Y$ such that $Y-K$ is irreducible and not homeomorphic to the solid torus.
-If two different Dehn surgeries on $K$ are purely cosmetic, then there is a homeomorphism of $Y-K$ which takes one slope to the other.
+Suppose $K$ is a knot in a closed oriented three-manifold $Y$ such that YKY-K is irreducible and not homeomorphic to the solid torus.
+If two different Dehn surgeries on $K$ are purely cosmetic, then there is a homeomorphism of YKY-K which takes one slope to the other.
 The existence of cosmetic surgeries has been ruled out in many cases for knots in $S^{3}$.
@@
 Much more is known when considering purely cosmetic surgeries.
-Hanselman proved that any pair of purely cosmetic Dehn surgeries on knots in $S^{3}$ must have slopes that are ±$\pm 2$ or ±$\pm 1/q$ for some integer $q$ [Han19].
+Hanselman proved that any pair of purely cosmetic Dehn surgeries on knots in $S^{3}$ must have slopes that are $\pm 2$ or $\pm 1/q$ for some integer $q$ [HAN19].
 Less is known for knots in arbitrary three-manifolds. Ni found an obstruction to cosmetic surgeries on knots in three-manifolds satisfying a certain condition on their Thurston norms [NI10].
@@
 Let $Y$ be a closed, oriented three-manifold, and $K\subset Y$ a null-homologous knot.
-We say that $K$ is blow-up unknotted if there exists a taut Seifert surface $F$ for $K$, some natural number $N$, and a properly-embedded smooth disk $G$ inside $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ with $\partial G=K\times\{0\}$ and $[G]=\iota_{*}[F]\in H_{2}(X,K\times\{0\})$ for $\iota:(Y,K)\to(X,K\times\{0\})$ the inclusion.
+We say that $K$ is blow-up bounding if there exists a taut Seifert surface $F$ for $K$, some natural number $N$, and a properly-embedded smooth disk $G$ inside $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ with $\partial G=K\times\{0\}$ and $G]=\iota_{*}[F]\in H_{2}(X,K\times\{0\})$ for $\iota:(Y,K)\to(X,K\times\{0\})$ the inclusion.
+We say that $K$ is blow-up unknotted if $K$ is obtained from the unknot in $Y$ by a sequence of crossing changes that change a positive crossing to a negative one.
+By Kirby calculus, $K$ is blow-up bounding if $K$ is blow-up unknotted.
 Remark.
-The property of $K$ being blow-up unknotted implies that $K$ is null homotopic.
-For example, by Kirby calculus, $K$ is a blow-up unknotted knot if $K$ is obtained from the unknot in $Y$ by a sequence of crossing changes that change a positive crossing to a negative one.
+The property of $K$ being blow-up bounding implies that $K$ is null homotopic.
 Theorem 1.4.
@@
 Theorem 1.5.
-Let $K\subset Y$ be a nontrivial blow-up unknotted knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $g(K)\geq 2$, then $Y_{1}(K)\ncong\pm Y$.
+Let $K\subset Y$ be a nontrivial blow-up bounding knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $K$ is a nontrivial blow-up unknotted knot, then $Y_{1}(K)\ncong\pm Y$.
 The proofs of these theorems rely on a mapping cone formula that relates the Heegaard Floer homology of Dehn surgery on null-homologous knots in arbitrary closed, oriented three-manifolds with the knot Floer chain complex.
@@
 Perhaps of independent interest, we provide a conceptually-simpler proof of the integer surgery formula for nontorsion $\operatorname{Spin}^{c}$ structures, which generalizes Ozsváth and Szabó’s and Ni’s results.
-Acknowledgements The author would like to thank Yi Ni for advising him in this project.
-2. Preliminaries on Heegaard Floer Homology
-In this section, we establish some notation and collect some results that will be needed in later sections.
+2 Preliminaries on Heegaard Floer Homology
+In this section, we establish some notation and collect some results, most of which are copied from [OS08], that will be needed in later sections.
 Let $W:Y_{1}\to Y_{2}$ be an oriented cobordism, with $Y_{1},Y_{2}$ connected. Given $\mathfrak{s}\in\operatorname{Spin}^{c}(W)$, there is a homomorphism
@@
 $\psi^{+}:MC(f_{2}^{+})\to CF^{+}(Y_{n}(K)).$
-Knot Floer homology associates to a knot $K$ in a closed, oriented three-manifold $Y$ a ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered ${\mathbb{Z}}[U]$-complex $C=CFK^{\infty}(Y,K)$.
-Write the two factors of the ${\mathbb{Z}}\oplus{\mathbb{Z}}$ filtration as $(i,j)$.
-Let $S$ be a region in the plane such that for all $(i,j)\in S$ and all $(i^{\prime},j^{\prime})\geq(i,j)$, we have that $(i^{\prime},j^{\prime})\in S$.
-Define $A_{s}^{+}=C\{\max(i,j-s)\geq 0\}$ and $B^{+}=C\{i\geq 0\}$ as in Ozsváth-Szabó.
+[…]
+Write the two factors of the $\mathbb{Z}$$\oplus{\mathbb{Z}}$ filtration as $(i,j)$. There is a $U$-action on $C$ that is an isomorphism of bifiltration (1,1)(-1,-1).
+We will often use the following fact, for example see [HW18, Reduction Lemma].
+Lemma 2.3.
+With coefficients in any field $\mathbb{F}$}, the knot Floer complex $(C,\partial)$ for any knot $K\subset Y$ is $({\mathbb{Z}}\oplus{\mathbb{Z}})$-filtered chain homotopy equivalent to a chain complex $(\overline{C},\overline{\partial})$ such that the restriction of the differential $\overline{\partial}$ to any $\overline{C}\{i,j\}$ is 0 and $\overline{C}\{0,j\}$ is isomorphic to $\widehat{HFK}(K,j)$.
+Knot Floer homology is known to detect the genus of the knot, and this is true for knots in general three-manifolds.
+Theorem 2.4.
+[NI10, Theorem 2.4] For any field $\mathbb{F}$}, if $K\subset Y$ is a knot of genus $g$, then
+$\widehat{HFK}(Y,K,s;{\mathbb{F}})=0$
+if $|s|>g$. Furthermore, $\widehat{HFK}(Y,K,g;{\mathbb{F}})\neq 0$.
+Define $A_{s}^{+}=C\{\max(i,j-s)\geq 0\}$ and $B^{+}=C\{i\geq 0\}$; these are chain complexes because the differential decreases both filtrations.
 There are two canonical chain maps $v_{s}^{+}:A_{s}^{+}\to B^{+}$ and $h_{s}^{+}:A_{s}^{+}\to B^{+}$, where $v_{s}^{+}$ is the vertical projection, and $h_{s}^{+}$ first projects $A_{s}^{+}$ to $C\{j\geq s\}$, then maps to $C\{j\geq 0\}$ via $U^{s}$, and finally maps to $B^{+}$ via a chain homotopy equivalence $C\{j\geq 0\}\to C\{i\geq 0\}$ [OS08].
+For all $s\in{\mathbb{Z}}$, there is a chain homotopy equivalence $\phi$ between $A_{s}^{+}$ and $A_{-s}^{+}$: first, the $U^{s}$ map is an isomorphism
+$A_{s}^{+}=C\{\max(i,j-s)\geq 0\}\to C\{\max(i,j-s)\geq-s\},$
+then the change of basepoint map that swaps the roles of $w$ and $z$ in the doubly-pointed Heegaard diagram is a chain homotopy equivalence
+$C\{\max(i,j-s)\geq-s\}\to C\{\max(i,j+s)\geq 0\}=A_{-s}^{+}.$
+Let $\phi$ be the composition of these two maps. It is clear from the definitions that the following lemma holds:
+Lemma 2.5.
+For all $s\in{\mathbb{Z}}$, $v_{s}^{+}$ and $h_{s}^{+}$ are chain homotopy equivalent to $h_{-s}^{+}\circ\phi$ and $v_{-s}^{+}\circ\phi$, respectively.
+Since all maps involved are $U$-equivariant, the above lemma also holds for the $CF^{\delta}$ version of knot Floer homology for any $\delta\geq 0$.
+Lemma 2.6.
+Let $g$ be the genus of $K$, then for all $\delta\geq 0$ and $s\geq g+\delta$, $v_{s}^{\delta}$, and $h_{-s}^{\delta}$ induce isomorphisms on homology, and $v_{-s-1}^{\delta}$ and $h_{s+1}^{\delta}$ are nullhomotopic.
+Proof.
+It is shown in the proof of the adjunction inequality in [OS04a, Theorem 5.1] that one can construct a special Heegaard diagram for $(Y,K)$ such that all generators of $\widehat{CFK}(Y,K)$ lie in Alexander grading greater than or equal to g-g.
+Thus, all generators for $A_{s}^{\delta}$ with $i\geq 0$ must have $j\geq-g$, from which it follows that $h_{-s}^{\delta}$ induces an isomorphism on homology and $v_{-s-1}^{\delta}$ is null-homotopic.
+By the $CF^{\delta}$ version of Lemma 2.5, we obtain the statements about $v_{s}^{\delta}$ and $h_{s+1}^{\delta}$.
+∎
 Let $K\subset Y$ be null-homologous.
-Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, and let $n\in{\mathbb{Z}}$.
-The space of $\operatorname{Spin}^{c}$ structures over $Y_{n}(K)$ which are $\operatorname{Spin}^{c}$-cobordant to $\mathfrak{t}$ over the cobordism $W_{n}^{\prime}(K)$ is identified with ${\mathbb{Z}}/n{\mathbb{Z}}$ by $\mathfrak{u}\mapsto i$, where $\mathfrak{s}$ is an extension of $\mathfrak{u}$ over $W_{n}^{\prime}(K)$ with the property that
-$\langle c_{1}(\mathfrak{s}),[\hat{F}]\rangle-n\equiv 2i\mod 2n.$
-For $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ denote the corresponding summand of
-$CF^{+}(Y_{n}(K))$. Let $A_{s,\mathfrak{t}}^{+},B_{\mathfrak{t}}^{+}$ denote the corresponding chain complexes associated to the knot filtration on $CF^{+}(Y,\mathfrak{t})$ [OS08].
+Let $F$ be a Seifert surface for $K$ and let $\hat{F}$ be the closed surface obtained by capping off $\partial F$ with a disk.
+Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, which need not have torsion first Chern class, and let $n\in{\mathbb{Z}}$.
+The space of $\operatorname{Spin}^{c}$ structures over $Y_{n}(K)$ which are $\operatorname{Spin}^{c}$-cobordant to $\mathfrak{t}$ over the natural two-handle cobordism $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ is identified with ${\mathbb{Z}}/n{\mathbb{Z}}$ by $\mathfrak{u}\mapsto i$, where $\mathfrak{s}$ is an extension of $\mathfrak{u}$ over $W_{n}^{\prime}(K)$ with the property that
+$\langle c_{1}(\mathfrak{s}),[\hat{F}]\rangle-n\equiv 2i\pmod{2n}$.
+For $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ denote the corresponding summand of $CF^{+}(Y_{n}(K))$.
+Let $A_{s,\mathfrak{t}}^{+},B_{\mathfrak{t}}^{+}$ denote the corresponding chain complexes associated to the knot filtration on $CF^{+}(Y,\mathfrak{t})$ [OS08].
 Let
-${\mathbb{A}}_{i,\mathfrak{t}}^{+}=\prod_{s\in{\mathbb{Z}}|s\equiv i\mod n}A_{s,\mathfrak{t}}^{+}\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{\mathfrak{t}}^{+}=\prod_{s\in{\mathbb{Z}}|s\equiv i\mod n}B_{\mathfrak{t}}^{+}$
+${\mathbb{A}}_{i,\mathfrak{t}}^{+}=\bigoplus_{s\in{\mathbb{Z}}|s\equiv i\pmod{n}}A_{s,\mathfrak{t}}^{+}\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{\mathfrak{t}}^{+}=\bigoplus_{s\in{\mathbb{Z}}|s\equiv i\pmod{n}}B_{\mathfrak{t}}^{+}$
 and let $D_{n,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{i,\mathfrak{t}}^{+}$ be the map
@@
 $b_{s}=h_{s-n}^{+}(a_{s-n})+v_{s}^{+}(a_{s}).$
-Let ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ denote the mapping cone of $D_{n,i,\mathfrak{t}}^{+}$ [OS08].
+Let ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ denote the mapping cone of $D_{n,i,\mathfrak{t}}^{+}$; we prove in the next section that this chain complex computes $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ [OS08].
 Taking the direct sum of the above chain complexes over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/n{\mathbb{Z}}$ gives $\mathbb{A}$$^{+}$, $\mathbb{B}$$^{+}$, and the map $D_{n}^{+}:{\mathbb{A}}^{+}\to{\mathbb{B}}^{+}$.
@@
 Let $K\subset Y$ be a null-homologous knot and let $F$ be a Seifert surface for $K$ in the homology class $\phi$.
-Given $n>0$ and $t\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ be the natural two-handle cobordism, and let $\widehat{F}_{n}\subset W_{n}^{\prime}(K)$ be the closed surface obtained by capping of $\partial F$ with a disk.
+Given $n>0$ and $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $\widehat{F}_{n}\subset W_{n}^{\prime}(K)$ be the closed surface obtained by capping off $\partial F$ with a disk.
 Let $\widehat{\phi}_{n}=[\widehat{F}_{n}]\in H_{2}(W_{n}^{\prime}(K))$.
-Let $k$ be an integer satisfying $k\equiv t\mod n$ and $|k|\leq n/2$.
-Let $\mathfrak{x}_{k},\mathfrak{y}_{k}\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ satisfy $\mathfrak{t}_{k}|Y=\mathfrak{n}_{k}|Y=\mathfrak{s}$ and
-$\langle c_{1}(\mathfrak{x}_{k}),\hat{\phi}_{n}\rangle=2k-n\hskip 14.22636pt\langle c_{1}(\mathfrak{y}_{k}),\hat{\phi}_{n}\rangle=2k+n,$
-and let $\mathfrak{s}_{t}=\mathfrak{t}_{k}|Y_{n}(K)=\mathfrak{n}_{k}|Y_{n}(K)$.
+Let $k$ be an integer satisfying $k\equiv i\pmod{n}$ and $|k|\leq n/2$.
+Let $\mathfrak{x}_{k},\mathfrak{y}_{k}\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ satisfy $\mathfrak{x}_{k}|Y=\mathfrak{y}_{k}|Y=\mathfrak{t}$ and
+$\langle c_{1}(\mathfrak{x}_{k}),\hat{\phi}_{n}\rangle=2k-n,\hskip 14.22636pt\langle c_{1}(\mathfrak{y}_{k}),\hat{\phi}_{n}\rangle=2k+n,$
+and let $\mathfrak{t}_{i}=\mathfrak{x}_{k}|Y_{n}(K)=\mathfrak{y}_{k}|Y_{n}(K)$.
 Theorem 2.7 ([OS08] Theorem 2.3).
-When $n\geq 2g(F)$, the chain complex
-$CF^{+}(Y_{n}(K),\mathfrak{s}_{t})$ is represented by the chain complex $A_{k}^{+}$, in the sense that there is an isomorphism
-$\Psi_{n}^{+}:CF^{+}(Y_{n}(K),\mathfrak{s}_{t})\to A_{k}^{+}.$
-Moreover, the following squares commute:
-{tikzcd}
-{tikzcd}
+When $n\geq 2g(F)$, the chain complex $CF^{+}(Y_{n}(K),\mathfrak{t}_{i})$ is represented by the chain complex $A_{k,\mathfrak{t}}^{+}$, in the sense that there is an isomorphism
+$\Psi_{n}^{+}:CF^{+}(Y_{n}(K),\mathfrak{t}_{i})\to A_{k,\mathfrak{t}}^{+}.$
+Moreover, the squares in the diagrams of Figure 1 commute.
+[…]
+[…]
+Figure 1: Commutative diagrams of the maps $v_{k}^{+}$ and $h_{k}^{+}$
+The $CF^{\delta}$ version of the theorem holds with the same proof.
 In the case of rational surgeries, let $p,q$ be a pair of relatively prime integers. For $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/p{\mathbb{Z}}$, denote
-${\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}=\prod_{r\in{\mathbb{Z}}}(r,A_{\lfloor(i+pr)/q\rfloor,\mathfrak{t}}^{+}(K))\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}=\prod_{r\in{\mathbb{Z}}}(r,B_{\mathfrak{t}}^{+}).$
+${\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}=\bigoplus_{r\in{\mathbb{Z}}}(r,A_{\lfloor(i+pr)/q\rfloor,\mathfrak{t}}^{+}(K))\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}=\bigoplus_{r\in{\mathbb{Z}}}(r,B_{\mathfrak{t}}^{+}).$
 Let $D_{p/q,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}$ be the map
@@
 $b_{r}=v_{\lfloor(i+pr)/q\rfloor}^{+}(a_{r})+h_{\lfloor(i+p(r-1))/q\rfloor}^{+}(a_{r-1}).$
-Let ${\mathbb{X}}_{p/q,i,\mathfrak{t}}^{+}$ denote the mapping cone of $D_{p/q,i,\mathfrak{t}}^{+}$ [OS10].
+Let ${\mathbb{X}}_{p/q,i,\mathfrak{t}}^{+}$ denote the mapping cone of $D_{p/q,i,\mathfrak{t}}^{+}$; in section 4, we prove that the homology of this chain complex is isomorphic to $HF^{+}(Y_{p/q}(K),i,\mathfrak{t})$ [OS10].
 As before, direct summing over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ and $i\in{\mathbb{Z}}/p{\mathbb{Z}}$ gives ${\mathbb{A}}_{p/q}^{+}$, ${\mathbb{B}}_{p/q}^{+}$, and $D_{p/q}^{+}:{\mathbb{A}}_{p/q}^{+}\to{\mathbb{B}}_{p/q}^{+}$.
@@
 In the case where $p/q$ is clear from the context, we will drop the subscript $p/q$ from the notation.
-3. Integer Surgeries
+3 Integer Surgeries
 In this section, we provide a proof of the mapping cone formula for integer surgeries in the case where the $\operatorname{Spin}^{c}$ structure has nontorsion first Chern class.
@@
 The long exact sequence in Theorem 2.2 decomposes over $\operatorname{Spin}^{c}$ structures on $Y$.
-Setting $m=nk$ for large $k$, it also decomposes over $\operatorname{Spin}^{c}$ structures on $Y_{n}(K)$, i.e.
+Setting $m=nk$ for large $k>0$, it also decomposes over $\operatorname{Spin}^{c}$ structures on $Y_{n}(K)$, i.e.
 given $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ where
-(3.1) $f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$
-$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k+1)}(K))$ $|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
-$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k+1)\equiv 2s\mod 2n(k+1),$
-$\text{and }s\equiv i\mod n\},$
-$CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ |\ s\equiv i\mod n}CF^{+}(Y_{n(k+1)}(K),s,\mathfrak{t}).$
-A strengthening of the proof of Proposition 3.5 in [Ni20] shows the following:
+$f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$ (3.1)
+$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k+1)}(K))\$$|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
+$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k+1)\equiv 2s\pmod{2n(k+1)}$,
+$\text{and }s\equiv i\pmod{n}\}$,
+$CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ |\ s\equiv i\pmod{n}}CF^{+}(Y_{n(k+1)}(K),s,\mathfrak{t}).$
+A strengthening of Proposition 3.5 in [NI20] shows the following:
 Proposition 3.2 ([NI20] Proposition 3.5).
-Let $p=\max\{k+g(F),-2k+g(F)\}$.
+Let $n>0,$ $n/2\leq k<n/2$, and $p=\max\{k+g(F),-2k+g(F)\}$.
 Suppose that $n\geq\max\{p,2g(F)\}$, $\mathfrak{s}\in\mathcal{X}_{k}\setminus\{\mathfrak{x}_{k},\mathfrak{y}_{k}\},$ then the map $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through the map
@@
 $U^{n-p}:\operatorname{HF}^{+}(Y_{n}(K),k,\mathfrak{t})\circlearrowright.$
-Proof.
-The proof is exactly the same as in [Ni20], except we note that
-$\frac{l(l+1)n}{2}-lk-g\geq n-p$
-whenever $l\neq 0,-1$. From here, it follows that $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through $U^{n-p}$. ∎
+This proposition is exactly the same as [NI20] Proposition 3.5 when $k\leq 0$, but when $k>0$, we improve the bound on the power of $U$ from n2kg(F)n-2k-g(F) to np=nkg(F)n-p=n-k-g(F).
+The proof of the improved bound is just checking that when $k>0$, the inequality
+$\frac{l(l+1)n}{2}-lk-g\geq n-2|k|-g$
+in Ni’s proof can be improved to
+$\frac{l(l+1)n}{2}-lk-g\geq n-k-g.$
+By the same proof, the proposition also holds for the $\delta$ version of Heegaard Floer homology for any $\delta\geq 0$.
 Theorem 3.3.
@@
 $D_{n,i,\mathfrak{t}}^{+}:{\mathbb{A}}_{i,\mathfrak{t}}^{+}\to{\mathbb{B}}_{i,\mathfrak{t}}^{+}$
-is identified with $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
+is quasi-isomorphic to $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
 Proof when $n>0$.
-Recall that $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ as in Equation 3.1.
+We first prove the theorem for the $CF^{\delta}$ version of knot Floer homology for any $\delta\geq 0$, that is, the mapping cone ${\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$ is quasi-isomorphic to $CF^{\delta}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
+Recall that $CF^{\delta}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{\delta})$ as in Equation 3.1.
 Since $c_{1}(\mathfrak{t})$ is nontorsion, $H_{*}(A_{i}^{+})$ is a finitely generated abelian group.
-So $U^{p}|H_{*}(A_{i}^{+})=0$ when $p$ is greater than a constant $C_{1}$ independent of $n,k$.
-By Theorem 2.3, this implies that $U^{p}|HF^{+}(Y_{n(k+1)}(K),i,\mathfrak{t})=0$ when $p>C_{1}$, $n(k+1)>2g$.
-Recall that we choose representatives
-$-\frac{n(k+1)}{2}\leq s<\frac{n(k+1)}{2}$
+So $U^{q}|H_{*}(A_{i}^{+})=0$ when $q$ is greater than a constant $C_{1}$ independent of $n,k$.
+Hence, $U^{q}|H_{*}(A_{i}^{\delta})=0$ for $q>C_{1}$ for the same constant $C_{1}$ independent of $n,k,$ and $\delta$.
+By the $CF^{\delta}$ version of Theorem 2.7, this implies that $U^{q}|HF^{\delta}(Y_{n(k+1)}(K),i,\mathfrak{t})=0$ when $q>C_{1}$ and $n(k+1)>2g$.
+Choose $k$ such that $n(k+1)>2(g+\delta)$.
+Consider a $\operatorname{Spin}^{c}$ structure $\mathfrak{s}_{s}\in\mathcal{X}_{i}$ for some $s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}$. Recall that we choose representatives
+$\frac{n(k+1)}{2}\leq s<\frac{n(k+1)}{2}$
 for classes in ${\mathbb{Z}}/n(k+1){\mathbb{Z}}$. Suppose $s\geq 0$. Then for $l=\max\{s+g,-2s+g\}$, we have
-$l\leq\frac{n(k+1)}{2}+g\leq n(k+1)$
-for large $k$. Proposition 3.2 implies that $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}=0$ when $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$.
-Now for $s<0$, let $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then by conjugation invariance, there is a chain equivalence
-$F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}\cong F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{+}.$
+$l\leq\frac{n(k+1)}{2}+g\leq n(k+1).$
+Proposition 3.2 implies that $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}=0$ when $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$.
+Now for $s<0$, let $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then by conjugation invariance, there is a chain homotopy equivalence
+$F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}\cong F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{\delta}.$
 If $\mathfrak{s}_{s}=\mathfrak{x}_{s}+r\operatorname{PD}[\widehat{F}$], then
 $\overline{\mathfrak{s}_{s}}=\overline{\mathfrak{x}}_{-s}-(r+1)\operatorname{PD}[\widehat{F}$],
-in particular, $\overline{\mathfrak{s}_{s}}$ is not one of the two special $\operatorname{Spin}^{c}$ structures and $-s>0$.
-By the above argument, $F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{+}=0$, hence also $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}=0$.
-The conclusion follows by Theorem 2.3. ∎
-3.1. Proof of Theorem for negative surgery coefficients
+in particular, $\overline{\mathfrak{s}_{s}}$ is not one of the two special $\operatorname{Spin}^{c}$ structures and s>0-s>0.
+By the above argument, $F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{\delta}=0$, hence also $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}=0$.
+Therefore, $MC(f_{2,i,\mathfrak{t}}^{\delta})$ is quasi-isomorphic to the mapping cone of
+$\sum_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}\left(f_{W_{n(k+1)}^{\prime}(K),\mathfrak{x}_{s}}^{\delta}+f_{W_{n(k+1)}^{\prime}(K),\mathfrak{y}_{s}}^{\delta}\right):CF^{\delta}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{\delta}(Y,\mathfrak{t}).$
+By Theorem 2.7, this chain complex is quasi-isomorphic to the mapping cone of
+$\sum_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}):\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}A_{s}^{\delta}\to\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}B^{\delta}.$
+We rewrite this as the truncated mapping cone
+$\sum_{-n(k+1)/2\leq s<n(k+1)/2\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}).$
+We chose $k$ large enough such that $\frac{n(k+1)}{2}>g+\delta$, so by Lemma 2.6, the above mapping cone is quasi-isomorphic to the mapping cone of
+$\sum_{s\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}):\bigoplus_{s\ :\ s\equiv i\pmod{n}}A_{s}^{\delta}\to\bigoplus_{s\ :\ s\equiv i\pmod{n}}B^{\delta},$
+which by definition is ${\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$.
+Therefore, we have the quasi-isomorphism
+$CF^{\delta}(Y_{n}(K),i,\mathfrak{t})\cong{\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$
+for all $\delta\geq 0$. Applying [OS08, Lemma 2.7], we conclude that
+$CF^{+}(Y_{n}(K),i,\mathfrak{t})\cong{\mathbb{X}}_{i,\mathfrak{t}}^{+}(n).$
+Proof of Theorem 3.3 for negative surgery coefficients.
 Continue to assume $n>0$, only we consider $Y_{-n}(K)$.
+Again, we first work with the $\delta$ version for $\delta\geq 0$.
 In this case, let $W^{\prime}_{n(k-1)}(K)$ be the cobordism from $Y_{n(k-1)}(K)$ to $Y$.
 For each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/n{\mathbb{Z}}$,
-$HF^{+}(Y_{-n}(K),i,\mathfrak{t})\cong H_{*}\left(MC\left(f_{2,i,\mathfrak{t}}^{+}:CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to\bigoplus^{nk}CF^{+}(Y,\mathfrak{t})\right)\right),$
+$HF^{\delta}(Y_{-n}(K),i,\mathfrak{t})\cong H_{*}\left(MC\left(f_{2,i,\mathfrak{t}}^{\delta}:CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to\bigoplus^{nk}CF^{\delta}(Y,\mathfrak{t})\right)\right),$
 where
-$f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k-1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$
-$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k-1)}(K))$ $|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
-$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k-1)\equiv 2s\mod 2n(k-1),$
-$\text{and }s\equiv i\mod n\},$
-$CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k-1){\mathbb{Z}}\ |\ s\equiv i\mod n}CF^{+}(Y_{n(k-1)}(K),s,\mathfrak{t}).$
-Consider the summand $CF^{+}(Y_{n(k-1)}(K),s,\mathfrak{t})$ for
-$-\frac{n(k-1)}{2}\leq s<\frac{n(k-1)}{2},\hskip 14.22636pts\equiv i\mod n.$
-Then, as above, if $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then $F_{W^{\prime}_{n(k-1)}(K),\mathfrak{s}_{s}}^{+}=0$.
-The conclusion again follows by Theorem 2.3.
-4. Rational Surgeries
+$f_{2,i,\mathfrak{t}}^{\delta}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k-1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}:CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to CF^{\delta}(Y,\mathfrak{t}),$
+$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k-1)}(K))\$$|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
+$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k-1)\equiv 2s\pmod{2n(k-1)}$,
+$\text{and }s\equiv i\pmod{n}\}$,
+$CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k-1){\mathbb{Z}}\ |\ s\equiv i\pmod{n}}CF^{\delta}(Y_{n(k-1)}(K),s,\mathfrak{t}).$
+Consider the summand $CF^{\delta}(Y_{n(k-1)}(K),s,\mathfrak{t})$ for
+$\frac{n(k-1)}{2}\leq s<\frac{n(k-1)}{2},\hskip 14.22636pts\equiv i\pmod{n}$.
+Then, as above, if $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then $F_{W^{\prime}_{n(k-1)}(K),\mathfrak{s}_{s}}^{\delta}=0$.
+The conclusion follows again by the same argument as in the positive surgery case.
+4 Rational Surgeries
 In this section, we sketch the proof of the rational surgery formula:
@@
 Then, for each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/p{\mathbb{Z}}$, there is an isomorphism of ${\mathbb{Z}}[U]$-modules
-$H_{*}(\mathbb{X}_{i,\mathfrak{t},p/q}^{+})\cong HF^{+}(Y_{p/q}(K),i,\mathfrak{t}).$
+$H_{*}(\mathbb{X}_{p/q,i,\mathfrak{t}}^{+})\cong HF^{+}(Y_{p/q}(K),i,\mathfrak{t}).$
 Ozsváth and Szabó stated and proved this theorem in the case where $\mathfrak{t}$ has torsion first Chern class [OS10].
@@
 Their argument still holds when the $\operatorname{Spin}^{c}$ structure is nontorsion, as long as grading information is ignored.
-By Kirby calculus, $Y_{p/q}(K)$ can be realized by surgery with coefficient $a$ inside the knot $K\#O_{q/r}\subset Y\#(-L(q,r))$, where $a$ is the greatest integer smaller than or equal to $p/q$, $a=\lfloor p/q\rfloor$, and $p/q=a+r/q$, and $O_{q/r}\subset-L(q,r)$ is the knot which is obtained by viewing one component of the the Hopf link as a knot inside the lens space $-L(q,r)$ thought of as $-q/r$ surgery on the other component of the Hopf link [OS10].
+Let $a=\lfloor p/q\rfloor$ and $p/q=a+r/q$.
+Consider the Hopf link in $S^{3}$ and perform q/r-q/r surgery on one component of the link.
+Let $O_{q/r}\subset-L(q,r)$ be the other component of the link viewed as a knot in the resulting lens space after surgery.
+By Kirby calculus, $Y_{p/q}(K)$ can be realized by surgery with coefficient $a$ inside the knot $K\#O_{q/r}\subset Y\#(-L(q,r))$ [OS10].
 Thus, one obtains the rational surgery formula by proving a formula for Morse surgeries on rationally null-homologous knots and then applying a Kunneth principle for connected sums.
-Analogously to the case of integer surgeries on null-homologous knots, for $K\subset Y$ a rationally null-homologous knot, $\lambda$ a framing on $K$, and $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, Ozsváth and Szabó define the chain map $D_{\mathfrak{s}}^{+}:{\mathbb{A}}_{\mathfrak{s}}^{+}(Y,\overline{K})\to{\mathbb{B}}_{\mathfrak{s}}^{+}(Y,\overline{K})$.
+Analogously to the case of integer surgeries on null-homologous knots, for $K\subset Y$ a rationally null-homologous knot with an orientation (we use $\overline{K}$ to denote the oriented knot), $\lambda$ a framing on $K$, and $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, define the chain map $D_{\mathfrak{s}}^{+}:{\mathbb{A}}_{\mathfrak{s}}^{+}(Y,\overline{K})\to{\mathbb{B}}_{\mathfrak{s}}^{+}(Y,\overline{K})$ as in [OS10] Theorem 6.1.
 Let ${\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$ be the mapping cone of $D_{\mathfrak{s}}^{+}$ [OS10].
-Theorem 4.2 (Morse surgery formula [OS10]).
+Theorem 4.2 (Morse surgery formula [OS10] Theorem 6.1).
 Let $K\subset Y$ be a rationally null-homologous knot, and let $\lambda$ be a framing on $K$. Let $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, then
-$HF^{+}(Y_{\lambda}(K),\mathfrak{s})\cong H_{*}({\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$
+$HF^{+}(Y_{\lambda}(K),\mathfrak{s})\cong H_{*}({\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda))$
 as ${\mathbb{Z}}[U]$-modules.
-We recall the definition of $U$-knots from [OS10].
+We recall the definition of $U$-knots from [OS10] Definition 5.2.
 Definition 4.3.
-A $U$-knot is a knot in a rational homology three-sphere $Y$ with the property that the induced filtration of $CF^{\infty}(Y,\mathfrak{s})$ is trivial.
-More precisely, given any $\xi\in\underline{\operatorname{Spin}^{c}}(Y,K)$, $CFK^{\infty}(Y,K,\xi)$ is chain homotopy equivalent via a relatively ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered homotopy to the chain complex $R$ which is a free, rank one ${\mathbb{Z}}[U,U^{-1}]$ module with the trivial differential.
-Proposition 4.4.
+[…]
+Proposition 4.4 ([OS10], Corollary 5.3).
 If $K_{2}\subset Y_{2}$ is a $U$-knot, then for each $\xi_{1}\in\underline{\operatorname{Spin}}^{c}(Y_{1},\overline{K}_{1})$ and $\mathfrak{s}_{2}\in\operatorname{Spin}^{c}(Y_{2})$, there is some $\xi_{2}\in\underline{\operatorname{Spin}^{c}}(Y_{2})$ representing $\mathfrak{s}_{2}$, with the property that
 $CFK^{\infty}(Y_{1},\overline{K_{1}},\xi_{1})\cong CFK^{\infty}(Y_{1}\#Y_{2},\overline{K_{1}}\#\overline{K_{2}},\xi_{1}\#\xi_{2})$
-as ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered chain complexes [OS10].
-Lemma 4.5.
-$O_{q/r}\subset-L(q,r)$ is a $U$-knot. There is an affine identification $\phi$ of the $\operatorname{Spin}^{c}$ structures on $\widehat{HFK}(L(q,r),O_{q,r})$ with ${\mathbb{Z}}$ such that
-$\widehat{HFK}(L(q,r),O_{q,r},\phi(i))\cong\begin{cases}{\mathbb{Z}}&\text{if $0\leq i\leq q-1$,}\\ 0&\text{otherwise.}\end{cases}$
-[OS10]
+[…]
+Lemma 4.5 ([OS10], Lemma 7.1).
+$O_{q/r}\subset-L(q,r)$ is a $U$-knot. There is an affine identification $\phi$ of the $\operatorname{Spin}^{c}$ structures on $\widehat{HFK}(L(q,r),O_{q/r})$ with $\mathbb{Z}$} such that
+$\widehat{HFK}(L(q,r),O_{q/r},\phi(i))\cong\begin{cases}{\mathbb{Z}}&\text{if $0\leq i\leq q-1$,}\\ 0&\text{otherwise.}\end{cases}$
 Let $K\subset Y$ be a null-homologous knot.
-Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, let $H^{2}(Y,K,\mathfrak{t})\cong{\mathbb{Z}}$ denote the subgroup of $H^{2}(Y,K)$ corresponding to $\mathfrak{t}$, and also $H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})\cong{\mathbb{Z}}$.
-Lemma 4.6.
+As in section 2.2 of [OS10], $H^{2}(Y,K)$ is defined to be $H^{2}(Y-\operatorname{nd}(K),\partial\operatorname{nd}(K))$ and is an affine space for $\underline{\operatorname{Spin}^{c}}(Y,K)$.
+There is a natural map $\underline{\operatorname{Spin}^{c}}(Y,K)\to\operatorname{Spin}^{c}(Y)$ which allows us to identify
+$\operatorname{Spin}^{c}(Y)\cong\frac{\underline{\operatorname{Spin}^{c}}(Y,K)}{{\mathbb{Z}}\cdot\operatorname{PD}[\mu]}.$
+Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, and under the above identification, let $H^{2}(Y,K,\mathfrak{t})\cong{\mathbb{Z}}$ denote the subgroup of $H^{2}(Y,K)$ corresponding to $\mathfrak{t}$, and also $H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})\cong{\mathbb{Z}}$.
+With this notation, the following more general version of [OS10, Lemma 7.2] holds with analogous proof.
+Lemma 4.6 ([OS10], Lemma 7.2).
 Let $K\subset Y$ be a null-homologous knot. Under the connected sum
 $(Y,K)\#(-L(q,r),O_{q,r})\to(Y\#-L(q,r),K\#O_{q,r}),$
-the following diagram commutes:
-{tikzcd}
-where $f(x,y)=qx+y$.
+the diagram in Figure 2 commutes where $f(x,y)=qx+y$.
 Moreover, under this correspondence, if $K_{\lambda}$ is the pushoff of $K$ with respect to the integral framing $a$, then $\operatorname{PD}[K_{\lambda}]$ represents the element $p\in{\mathbb{Z}}\cong H^{2}(Y\#-L(q,r),K\#O_{q,r})$, where $p/q=a+r/q$.
+$\mathbb{Z}\oplus\mathbb{Z}$$\mathbb{Z}$$H^{2}(Y,K,\mathfrak{t})\oplus H^{2}(-L(q,r),O_{q,r})$$H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})$$\scriptstyle{f}\scriptstyle{\cong}\scriptstyle{\cong}$
+Figure 2:
 Proof.
@@
 With these ingredients, the proof of Theorem 4.1 follows as in [OS10].
-5. Cosmetic Surgeries
+5 Cosmetic Surgeries
 In this section, we prove our main results.
@@
 We state some lemmas that will be useful in calculating the homology of the mapping cones for Dehn surgery.
-The following lemma was proven in [Ni20].
+The following lemma was proven in [NI20] Equations (5) and (6).
 Lemma 5.1.
-If $s\leq t$, then $*\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{t})_{*}$ and $*\operatorname{Im}(\widehat{h}_{s})_{*}\supset\operatorname{Im}(\widehat{h}_{t})_{*}$.
+If $s\leq t$, then $\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{t})_{*}$ and $\operatorname{Im}(\widehat{h}_{s})_{*}\supset\operatorname{Im}(\widehat{h}_{t})_{*}$.
+As a consequence of the hat version of Lemma 2.5, $\operatorname{rank}H_{*}(\widehat{A}_{s})=\operatorname{rank}H_{*}(\widehat{A}_{-s})$ and $\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}(\widehat{h}_{-s})_{*}$ for all $s$.
 Lemma 5.2.
-There is an equivalence between $\widehat{v}_{s}$ and $\widehat{h}_{-s}$ for all $s$.
-This equivalence is induced by changing the roles of the basepoints $w$ and $z$ in the doubly-pointed Heegaard diagram.
-Lemma 5.3.
-Let $g(K)$ be the genus of $K$, then for all $s\geq g(K)$, $\widehat{v}_{s}$, $\widehat{h}_{-s}$ induce isomorphisms on homology.
-For all $s\geq g(K)+1$, $\widehat{v}_{-s}$ and $\widehat{h}_{s}$ are nullhomotopic.
+Suppose for some $s\in{\mathbb{Z}}$ that $\widehat{v}_{s}$ or $\widehat{h}_{s}$ induces an isomorphism on homology.
+Then $v_{s}^{+}$ or $h_{s}^{+}$ also induces an isomorphism on homology, respectively.
+Proof.
+We prove this lemma for $\widehat{v}_{s}$, and the case for $\widehat{h}_{s}$ is proven similarly.
+We will show by induction that $(v_{s}^{\delta})_{*}$ is an isomorphism for all $\delta\geq 0$, which by Lemma 2.7 in [OS08] implies that $(v_{g-1}^{+})_{*}$ is an isomorphism.
+Suppose that $(v_{s}^{\delta})_{*}$ is an isomorphism for some $\delta\geq 0$ (the base case $\delta=0$ is $v_{s}^{0}=\widehat{v}_{s}$).
+Clearly, the maps $\widehat{v}_{s}$, $v_{s}^{\delta+1}$, and $v_{s}^{\delta}$ fit into the commutative diagram of short exact sequences in Figure 3.
+Therefore, the diagram in Figure 4 commutes and the rows are exact, so by the five lemma, $(v_{s}^{\delta+1})_{*}$ is also an isomorphism.
+∎
+[…]
+Figure 3:
+$H_{*}(A_{s}^{\delta})$$H_{*}(\widehat{A}_{s})$$H_{*}(A_{s}^{\delta+1})$$H_{*}(A_{s}^{\delta})$$H_{*}(\widehat{A}_{s})$$H_{*}(B^{\delta})$${H_{*}(\widehat{B})}$$H_{*}(B^{\delta+1})$$H_{*}(B^{\delta})$${H_{*}(\widehat{B})}\scriptstyle{\partial}$$\scriptstyle{(v_{s}^{\delta})_{*}}$$\scriptstyle{i_{*}}$$\scriptstyle{(\widehat{v}_{s})_{*}}$$\scriptstyle{U_{*}}$$\scriptstyle{(v_{s}^{\delta+1})_{*}}\scriptstyle{\partial}$$\scriptstyle{(v_{s}^{\delta})_{*}}$$\scriptstyle{(\widehat{v}_{s})_{*}}\scriptstyle{\partial}$$\scriptstyle{i_{*}}$$\scriptstyle{U_{*}}\scriptstyle{\partial}$
+Figure 4:
 Lemma 5.3 ([NI20] Lemma 4.1).
@@
 Let $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ for some $N$, and let $\iota:(Y,K)\to(X,K\times\{0\})$ be the inclusion map.
-Let $G\subset X$ be a properly-embedded smooth connected surface with $\partial G=K\times\{0\}$, $[G]=\iota_{*}[F]\in H^{2}(X,K\times\{0\})$, and $g(G)<g(F)$.
+Let $G\subset X$ be a properly-embedded smooth connected surface with $\partial G=K\times\{0\}$, $G]=\iota_{*}[F]\in H^{2}(X,K\times\{0\})$, and $g(G)<g(F)$.
 Then
@@
 $(v_{0}^{+})_{*}=(h_{0}^{+})_{*}$
-Again, call a knot where such $(X,G)$ for $G$ a disk exists a blow-up unknotted knot. We also consider the unknot itself to be a blow-up unknotted knot.
+The same is true for the hat version of knot Floer homology by essentially the same proof.
+Again, call a knot where such $(X,G)$ for $G$ a disk exists a blow-up bounding knot. We also consider the unknot itself to be a blow-up bounding knot.
+The following argument is due to Ni:
+Lemma 5.4.
+[…]
+Proof.
+Denote by $O$ the unknot in $Y$.
+There is a sequence of $N$ crossing changes changing a positive crossing to a negative one that changes $O$ to the knot $K$.
+Let $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ and $W_{n}^{\prime}(O):Y_{n}(O)\to Y$ denote the natural two-handle cobordisms.
+Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ on $Y$, let $\mathfrak{x}_{0}(O)\in\operatorname{Spin}^{c}(W_{n}^{\prime}(O))$ and $\mathfrak{x}_{0}(K)\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ as in Section 2 restrict to the $\operatorname{Spin}^{c}$ strctures $\mathfrak{x}_{0}(O)|Y=\mathfrak{x}_{0}(K)|Y=\mathfrak{t}$ and $\mathfrak{t}_{0}(O)=\mathfrak{x}_{0}(O)|Y_{n}(O)$, $\mathfrak{t}_{0}(K)=\mathfrak{x}_{0}(K)|Y_{n}(K)$.
+Let $\mathfrak{S}$ be the $\operatorname{Spin}^{c}$ structure on $W_{n}^{\prime}(O)\#N\overline{{\mathbb{C}}P^{2}}$ given by the connected sum of $\mathfrak{x}_{0}(O)$ with $N$ copies of $\mathfrak{s}_{1}\in\operatorname{Spin}^{c}(\overline{{\mathbb{C}}P^{2}})$, where $\mathfrak{s}_{1}$ satisfies that $c_{1}(\mathfrak{s}_{1})$ generates $H^{2}(\overline{{\mathbb{C}}P^{2}})$.
+By Kirby calculus, the Kirby diagram of the $n$-framed unknot in $Y$ blown up $N$ times is equivalent to the diagram of the $n$-framed knot $K$ in $Y$.
+Thus, the cobordism $W_{n}^{\prime}(O)\#N\overline{{\mathbb{C}}P^{2}}$ is the composition of two cobordisms
+$Y_{n}(O)\xrightarrow{X_{n}(K)}Y_{n}(K)\xrightarrow{W_{n}^{\prime}(K)}Y$
+for some cobordism $X_{n}(K)$.
+The $\operatorname{Spin}^{c}$ structure $\mathfrak{S}$ restricts to $\mathfrak{t}_{0}(O)$ on $Y_{n}(O)$, $\mathfrak{t}_{0}(K)$ on $Y_{n}(K)$, and $\mathfrak{t}$ on $Y$.
+Therefore, $X_{n}(K)$ induces a cobordism map
+$\Phi=\widehat{F}_{X_{n}(K),\mathfrak{S}|X_{n}(K)}:\widehat{HF}(Y_{n}(O),\mathfrak{t}_{0}(O))\to\widehat{HF}(Y_{n}(K),\mathfrak{t}_{0}(K))$
+that by Theorem 2.7 fits into the commutative diagram in Figure 5.
+By a sample calculation, $(\widehat{v}_{0}(O))_{*}$ is surjective.
+Therefore $(\widehat{v}_{0}(K))_{*}$ must also be surjective.
+Direct summing over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ proves the statement.
+$\widehat{HF}(Y_{n}(O),\mathfrak{t}_{0}(O))$$\widehat{HF}(Y_{n}(K),\mathfrak{t}_{0}(K))$$H_{*}(\widehat{A}_{0,\mathfrak{t}}(O))$$H_{*}(\widehat{A}_{0,\mathfrak{t}}(K))$$H_{*}(\widehat{B}_{\mathfrak{t}}(O))$$H_{*}(\widehat{B}_{\mathfrak{t}}(K))$$\scriptstyle{\Phi}\scriptstyle{\cong}\scriptstyle{\cong}$$\scriptstyle{(\widehat{v}_{0}(O))_{*}}$$\scriptstyle{(\widehat{v}_{0}(K))_{*}}$$\scriptstyle{=}$
+Figure 5:
 The following result and its proof are analogous to [OS10] Proposition 9.7.
@@
 In particular, $K$ is the unknot if and only if $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$.
-The proof makes use of the $CF^{\delta}$ variant of Heegaard Floer homology defined in [OS08].
-For any nonnegative integer $\delta$, Denote by $CF^{\delta}(Y)$ the subcomplex of $CF^{+}(Y)$ which is the kernel of multiplication by $U^{\delta+1}$.
-Similarly, let $A_{s}^{\delta}$, $B^{\delta}$ denote the subcomplexes of $A_{s}^{+}$, $B^{+}$, respectively, which are the kernel of multiplication by $U^{\delta+1}$.
 Proof.
@@
 Consider the short exact sequence
-(5.6) $0\to C\{i<0,j\geq g-1\}\to C\{i\geq 0\text{ or }j\geq g-1\}\xrightarrow{v_{g-1}^{+}}C\{i\geq 0\}\to 0$
-The middle term is $A_{g-1}^{+}$, and $C\{i\geq 0\}=B^{+}$. Also,
-$H_{*}(C\{i<0,j\geq g-1\})\cong H_{*}(C\{(-1,g-1)\}\cong H_{*}(C\{0,g\})=\widehat{HFK}(K,g)\neq 0.$
-Therefore, $(v_{g-1}^{+})_{*}$ is not an isomorphism.
-Suppose for the sake of contradiction that $(\widehat{v}_{g-1})_{*}$ is an isomorphism.
-Suppose also that $(v_{g-1}^{\delta})_{*}$ is an isomorphism for some $\delta\geq 0$.
-By [OS04a] Theorem 4.4, the diagram in Figure 1 commutes and the rows are exact, so by the five lemma, $(v_{g-1}^{\delta+1})_{*}$ is also an isomorphism.
-By induction, $(v_{g-1}^{\delta})_{*}$ is an isomorphism for all $\delta$.
-Thus, by Lemma 2.7 in [OS08], $(v_{g-1}^{+})_{*}$ is an isomorphism, which is a contradiction.
-{tikzcd}
-Figure 1.
+$0\to C\{i<0,j\geq g-1\}\to C\{i\geq 0\text{ or }j\geq g-1\}\xrightarrow{v_{g-1}^{+}}C\{i\geq 0\}\to 0.$ (5.6)
+The middle term is $A_{g-1}^{+}$, and $C\{i\geq 0\}=B^{+}$.
+The assumption that $C$ is reduced together with the $U$-isomorphism and the fact that knot Floer homology detects genus implies that $C\{i,j\}$ is nonzero only if $g\leq i-j\leq g$.
+Therefore,
+$H_{*}(C\{i<0,j\geq g-1\})\cong H_{*}(C\{(-1,g-1)\}),$
+which by the $U$-isomorphism is isomorphic to
+$H_{*}(C\{0,g\})=\widehat{HFK}(K,g),$
+and $\widehat{HFK}(K,g)\neq 0$ since knot Floer homology detects genus. Therefore, $(v_{g-1}^{+})_{*}$ is not an isomorphism. By Lemma 5.2, $(\widehat{v}_{g-1})_{*}$ is also not an isomorphism.
+We abuse notation to let $\widehat{A}_{j/q}$ denote $\widehat{A}_{\lfloor j/q\rfloor}$, and also $\widehat{B}_{j/q}$, $\widehat{v}_{j/q}$, $\widehat{h}_{j/q}$ similarly.
 Lemma 5.7.
-Let $K\subset Y$ be a blow-up unknotted knot, $p,q>0$ relatively prime, and let $x\in H_{*}(\widehat{A}_{j/q})$ for some $j$.
-If $\lfloor j/q\rfloor\geq 0$, there exists $\alpha\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ such that $(\widehat{D}_{p/q})_{*}(x+\alpha)=(\widehat{v}_{j/q})_{*}(x)$ (where we think of $H_{*}(\widehat{B}_{j/q})$ as including into $H_{*}(\widehat{\mathbb{B}})$).
-Similarly, if $\lfloor j/q\rfloor\leq 0$, there exists $\alpha\in H_{*}(\oplus_{i\leq j-p}\widehat{A}_{i/q})$ such that $(\widehat{D}_{p/q})_{*}(x+\alpha)=(\widehat{h}_{j/q})_{*}(x)$.
+[…]
 Proof.
-For $s\geq 0$, we have $*\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s+1})_{*}\subset\dots$, and for $s\leq 0$, $*\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s-1})_{*}\subset\dots$.
+Since $K$ is blow-up bounding, by Lemma 5.3, for $s\geq 0$, we have $\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s+1})_{*}\subset\dots$, and for $s\leq 0$, $\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s-1})_{*}\subset\dots$.
 Thus, in the first case, there exists $y_{1}\in H_{*}(\widehat{A}_{(j+p)/q})$ such that $(\widehat{h}_{j/q})_{*}(x)=(\widehat{v}_{(j+p)/q})_{*}(y_{1})$.
@@
 By Lemma 2.6, this process eventually terminates and gives an element $\alpha=\sum_{n\geq 1}y_{n}\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ with the desired condition.
-Similarly, in the case where $\lfloor j/q\rfloor\leq 0$, since $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\subset\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, we can inductively add cancellation terms to the left.
+Similarly, in the case where $\lfloor j/q\rfloor\leq 0$, since $\operatorname{Im}(\widehat{v}_{j/q})_{*}\subset\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, we can inductively add cancellation terms to the left.
 ∎
-Using the mapping cone formula and the above lemmas, in the case where $K$ is blow-up unknotted, we may now calculate the rank of $\widehat{HF}(Y_{r}(K))$ for any positive rational number $r$.
+Using the mapping cone formula and the above lemmas, in the case where $K$ is blow-up bounding, we may now calculate the rank of $\widehat{HF}(Y_{r}(K))$ for any positive rational number $r$.
 Lemma 5.8.
-Let $K\subset Y$ be a blow-up unknotted knot, let $p,q>0$ coprime. Let $b=\operatorname{rank}(\widehat{B})=\operatorname{rank}\widehat{HF}(Y)$. Then
-(5.9) $\begin{split}\operatorname{rank}\widehat{HF}(Y_{p/q}(K))&=q(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &+2q\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})+2t_{K}^{p/q}-pb,\end{split}$
+Let $K\subset Y$ be a blow-up bounding knot, let $p,q>0$ coprime. Let $b=\operatorname{rank}(H_{*}(\widehat{B}))=\operatorname{rank}\widehat{HF}(Y)$. Then
+$\begin{split}\operatorname{rank}\widehat{HF}(Y_{p/q}(K))&=q(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &+2q\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})+2t_{K}^{p/q}-pb,\end{split}$ (5.9)
 where
-(5.10) $t_{K}^{p/q}$ $=\sum_{j=0}^{p-1}\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right).$
+$t_{K}^{p/q}$ $=\sum_{j=0}^{p-1}\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right).$ (5.10)
 Note that if $Y=S^{3}$, Equation 5.9 agrees with Equation 40 in [OS10]. Indeed, define $\nu(K)$ as in [OS10]. If $\nu(K)>0$, then
@@
 Proof.
-By Theorem 4.1, $\widehat{HF}(Y_{p/q}(K))\cong H_{*}(\widehat{\mathbb{X}}_{p/q})$, the homology of the mapping cone of $\widehat{D}_{p/q}$ from $\widehat{{\mathbb{A}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{A}}_{i}$ to $\widehat{{\mathbb{B}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{B}}_{i}$.
-We abuse notation to let $\widehat{A}_{j/q}$ denote $\widehat{A}_{\lfloor j/q\rfloor}$, and also $\widehat{B}_{j/q}$, $\widehat{v}_{j/q}$, $\widehat{h}_{j/q}$ similarly.
+By Theorem 4.1, $\widehat{HF}(Y_{p/q}(K))\cong H_{*}(\widehat{\mathbb{X}}_{p/q})$, the homology of the mapping cone of $\widehat{D}_{p/q}$ from $\widehat{{\mathbb{A}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{A}}_{i}$ to $\widehat{{\mathbb{B}}}=\oplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\widehat{\mathbb{B}}_{i}$ (we have omitted the slope $p/q$ from the notation and summed over all $\operatorname{Spin}^{c}$ structures on $Y$).
 So
-$\widehat{\mathbb{A}}=\bigoplus_{i\in{\mathbb{Z}}/p{\mathbb{Z}}}\prod_{r\in{\mathbb{Z}}}\widehat{A}_{(i+pr)/q}=\prod_{j\in{\mathbb{Z}}}\widehat{A}_{j/q}$
+$\widehat{\mathbb{A}}=\bigoplus_{i=0}^{p-1}\prod_{r\in{\mathbb{Z}}}\widehat{A}_{(i+pr)/q}=\prod_{j\in{\mathbb{Z}}}\widehat{A}_{j/q}$
 has $q$ copies of $\widehat{A}_{s}$ for each $s\in{\mathbb{Z}}$, and under the map $\widehat{v}$, each $\widehat{A}_{j/q}$ corresponds to a copy of $\widehat{B}$ in $\widehat{\mathbb{B}}$, which we shall denote $\widehat{B}_{j/q}$.
 Then $\widehat{h}_{j/q}$ maps $\widehat{A}_{j/q}$ to $\widehat{B}_{(j+p)/q}$.
-Thus, for each $s\in{\mathbb{Z}}$, there are $p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s-1}$ and $q-p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s}$.
+Thus, for each $s\in{\mathbb{Z}}$, there are $p$ copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s-1}$ and qpq-p copies of $\widehat{B}_{s}$ in the codomain of $\widehat{h}_{s}$.
 The finiteness of $H_{*}(\widehat{\mathbb{X}}_{p/q})$ is guaranteed by Lemma 2.6; in fact, it implies that $\widehat{\mathbb{X}}_{p/q}$ is isomorphic to a truncated version of the mapping cone
@@
 $\widehat{\mathbb{A}}_{p/q}^{c}=\bigoplus_{j=-qc+1}^{qc-1}\widehat{A}_{j/q}=\bigoplus_{s=-c+1}^{c-1}\bigoplus_{i=0}^{q-1}\widehat{A}_{s},\hskip 14.22636pt\widehat{\mathbb{B}}_{p/q}^{c}=\bigoplus_{j=-qc+p+1}^{qc-1}\widehat{B}_{j/q},$
-and $\widehat{D}_{p/q}^{c}$ is the restriction of $\widehat{D}_{p/q}$ to $\widehat{A}_{p/q}^{c}$, with the exception that for any $x\in\widehat{A}_{j/q}$ with $-qc+1\leq j\leq-qc+p$, $\widehat{D}_{p/q}^{c}(x)=\widehat{h}_{j/q}(x)$, and for any $x\in\widehat{A}_{j/q}$ with $qc-p\leq j\leq qc-1$, $\widehat{D}_{p/q}^{c}(x)=\widehat{v}_{j/q}(x)$.
+and $\widehat{D}_{p/q}^{c}$ is the restriction of $\widehat{D}_{p/q}$ to $\widehat{A}_{p/q}^{c}$, with the exception that for any $x\in\widehat{A}_{j/q}$ with $qc+1\leq j\leq-qc+p$, $\widehat{D}_{p/q}^{c}(x)=\widehat{h}_{j/q}(x)$, and for any $x\in\widehat{A}_{j/q}$ with $qc-p\leq j\leq qc-1$, $\widehat{D}_{p/q}^{c}(x)=\widehat{v}_{j/q}(x)$.
 Note that it suffices to take $c\geq g+p/q+1$, and
 $\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})=q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s}),\hskip 14.22636pt\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})=qb(2c-1)-pb.$
-We construct a basis for $*\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}$. For $j\geq 0$, let
-$*\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$
-be a basis of $*\ker(\widehat{v}_{j/q})_{*}$.
-For each such $x_{j/q}^{i}$, choose an $\alpha_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\alpha_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$.
+We construct a basis for $\ker(\widehat{D}_{p/q})_{*}$ of the entire mapping cone, which has the same rank as $\ker(\widehat{D}_{p/q}^{c})_{*}$ for large $c$ by Lemma 2.6. For $j\geq 0$, let
+$\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$
+be a basis of $\ker(\widehat{v}_{j/q})_{*}$.
+For each such $x_{j/q}^{i}$, choose an $\alpha_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\alpha_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$.
 For $j<0$, let
-$*\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{h}_{j/q})_{*}}$
-be a basis of $*\ker(\widehat{h}_{j/q})_{*}$.
-For each such $x_{j/q}^{i}$, choose a $\beta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\beta_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$.
-Lastly, for $10\leq j\leq p-$ 1, let
+$\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{h}_{j/q})_{*}}$
+be a basis of $\ker(\widehat{h}_{j/q})_{*}$.
+For each such $x_{j/q}^{i}$, choose a $\beta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{x}_{j/q}^{i}=x_{j/q}^{i}+\beta_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$.
+Lastly, for $0\leq j\leq p-1$, let
 $\{y_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\left(\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}\right)}$
-be a basis of the subspace of $H_{*}(\widehat{A}_{j/q})$ that is mapped isomorphically to $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$ under $(\widehat{v}_{j/q})_{*}$.
+be a basis of a subspace of $H_{*}(\widehat{A}_{j/q})$ that is mapped isomorphically to $\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$ under $(\widehat{v}_{j/q})_{*}$.
 Choose $z_{j/q}^{i}\in H_{*}(\widehat{A}_{(j-p)/q})$ such that $(\widehat{v}_{j/q})_{*}(y_{j/q}^{i})=(\widehat{h}_{(j-p)/q})_{*}(z_{j/q}^{i})$.
-Choose $\gamma_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ and $\delta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{y}_{j/q}^{i}=y_{j/q}^{i}+z_{j/q}^{i}+\gamma_{j/q}^{i}+\delta_{j/q}^{i}$ is in $*\ker(\widehat{D}_{p/q})_{*}$.
+Choose $\gamma_{j/q}^{i}\in H_{*}(\oplus_{k\geq j+p}\widehat{A}_{k/q})$ and $\delta_{j/q}^{i}\in H_{*}(\oplus_{k\leq j-p}\widehat{A}_{k/q})$ according to Lemma 5.7 such that $\widetilde{y}_{j/q}^{i}=y_{j/q}^{i}+z_{j/q}^{i}+\gamma_{j/q}^{i}+\delta_{j/q}^{i}$ is in $\ker(\widehat{D}_{p/q})_{*}$.
 The set
@@
 is linearly independent, and it is finite by Lemma 2.6.
-Moreover, it spans $*\ker(\widehat{D}_{p/q})_{*}$: let $*z\in\ker(\widehat{D}_{p/q})_{*}$ be nonzero, and suppose first that $z\in H_{*}(\oplus_{k\geq 0}\widehat{A}_{k/q})$.
+Moreover, it spans $\ker(\widehat{D}_{p/q})_{*}$: let $z\in\ker(\widehat{D}_{p/q})_{*}$ be nonzero, and suppose first that $z\in H_{*}(\oplus_{k\geq 0}\widehat{A}_{k/q})$.
 Let $z_{j/q}$ denote the term of $z$ in $H_{*}(\widehat{A}_{j/q})$, and let
 $j=\min\{k\geq 0\ |\ z_{k/q}\neq 0\}.$
-Then $z\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, so $*z_{j/q}\in\operatorname{Span}\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$: otherwise if $*z_{j/q}\notin\ker(\widehat{v}_{j/q})_{*}$, then $*z\notin\ker(\widehat{D}_{p/q})_{*}$ since $z\in H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$.
-Thus,
+Then $z\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, so $z_{j/q}\in\operatorname{Span}\{x_{j/q}^{i}\}_{i=1}^{\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}}$: otherwise if $z_{j/q}\notin\ker(\widehat{v}_{j/q})_{*}$, then since $z\in H_{*}(\oplus_{k\geq j}\widehat{A}_{k/q})$, we have $z_{(j-p)/q}=0$ so
+$(\widehat{v}_{j/q})_{*}(z_{j/q})+(\widehat{h}_{(j-p)/q})_{*}(z_{(j-p)/q})\neq 0\in H_{*}(\widehat{B}_{j/q})$
+which means $z\notin\ker(\widehat{D}_{p/q})_{*}$. Thus, we can subtract some combination of the basis elements $\widetilde{x}_{j/q}^{i}$ to eliminate the $H_{*}(\widehat{A}_{j/q})$ component of $z$:
 $z-\sum_{i\in J}\widetilde{x}_{j/q}^{i}\in\ker(\widehat{D}_{p/q})_{*}\cap H_{*}(\oplus_{k>j}\widehat{A}_{k/q})$
-[…]
+for some indexing set $J\subset\{1,\dots,\operatorname{rank}\ker(\widehat{v}_{j/q})_{*}\}$. Inductively, we can then write
 $z=\sum_{k\in I}\sum_{i_{k}\in J_{k}}\widetilde{x}_{k/q}^{i_{k}}$
-[…]
+for some finite indexing sets $I\subset\{k\geq j\}$, $J_{k}\subset\{1,\dots,\operatorname{rank}\ker(\widehat{v}_{k/q})_{*}\}$.
+Similarly, if $z\in H_{*}(\oplus_{k<0}\widehat{A}_{k/q})$, then we can write
 $z=\sum_{k\in I}\sum_{i_{k}\in J_{k}}\widetilde{x}_{k/q}^{i_{k}}$
-[…]
-Now for the general case $*z\in\ker(\widehat{D}_{p/q})_{*}$, we claim that for each $10\leq j\leq p-$ 1,
+for some finite indexing set $I\subset\{k\leq j\}$, $J_{k}\subset\{1,\dots,\operatorname{rank}\ker(\widehat{h}_{k/q})_{*}\}$.
+Now for the general case $z\in\ker(\widehat{D}_{p/q})_{*}$, we claim that for each $0\leq j\leq p-1$,
 $z_{j/q}\in\operatorname{Span}\left(\{y_{j/q}^{i}\}\cup\{x_{j/q}^{i}\}\right).$
-Indeed, suppose this is not true for some $10\leq j\leq p-$ 1, then $*(\widehat{v}_{j/q})_{*}(z_{j/q})\notin\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, so $(\widehat{D}_{p/q})_{*}(z)\neq 0$, a contradiction.
+Indeed, suppose this is not true for some $0\leq j\leq p-1$, then $(\widehat{v}_{j/q})_{*}(z_{j/q})\notin\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, so $(\widehat{D}_{p/q})_{*}(z)\neq 0$, a contradiction.
 Thus,
@@
 By the above argument, $z$ is in the span of $S$.
-Therefore, $S$ is a basis for $*\ker(\widehat{D}_{p/q})_{*}$, so since $*\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}\ker(\widehat{h}_{-s})_{*}$,
+Therefore, $S$ is a basis for $\ker(\widehat{D}_{p/q})_{*}$, so since $\operatorname{rank}\ker(\widehat{v}_{s})_{*}=\operatorname{rank}\ker(\widehat{h}_{-s})_{*}$ by the hat version of Lemma 2.5 and these are 0 when $s\geq g$ by Lemma 2.6,
 $\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}=q\operatorname{rank}\ker(\widehat{v}_{0})_{*}+2q\sum_{s=1}^{g-1}\operatorname{rank}\ker(\widehat{v}_{s})_{*}+t_{K}^{p/q}.$
-Since $*\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})-\operatorname{rank}\ker(\widehat{v}_{s})_{*}$, $\operatorname{rank}H_{*}(\widehat{A}_{s})=b$ for all $|s|\geq g$, and $*\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}=\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})-\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})+\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*}$ for some large $c$, it follows that
-$*\displaystyle\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$ $=b(2c-1)q-pb-q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s})$
+Since $\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{s})-\operatorname{rank}\ker(\widehat{v}_{s})_{*}$ and $\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}=\operatorname{rank}H_{*}(\widehat{\mathbb{B}}_{p/q}^{c})-\operatorname{rank}H_{*}(\widehat{\mathbb{A}}_{p/q}^{c})+\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*}$ for some large $c$, it follows that
+$\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$ $=b(2c-1)q-pb-q\sum_{s=-c+1}^{c-1}\operatorname{rank}H_{*}(\widehat{A}_{s})$
 $\hskip 14.22636pt+q\operatorname{rank}\ker(\widehat{v}_{0})_{*}+2q\sum_{s=1}^{g-1}\operatorname{rank}\ker(\widehat{v}_{s})_{*}+t_{K}^{p/q}$
@@
 $=q(b-\operatorname{rank}(\widehat{v}_{0})_{*})+2q\sum_{s=1}^{g-1}(b-\operatorname{rank}(\widehat{v}_{s})_{*})+t_{K}^{p/q}-pb,$
-and therefore Equation 5.9 holds.
+where the second equality uses that $\operatorname{rank}H_{*}(\widehat{A}_{s})=b$ for all $|s|\geq g$ by Lemmas 2.5 and  2.6. Therefore, because
+$\operatorname{rank}(H_{*}(\widehat{\mathbb{X}}_{p/q}))=\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}+\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q})_{*}$
+and
+$\operatorname{rank}\ker(\widehat{D}_{p/q})_{*}=\operatorname{rank}\ker(\widehat{D}_{p/q}^{c})_{*},\hskip 14.22636pt\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q})_{*}=\operatorname{rank}\operatorname{coker}(\widehat{D}_{p/q}^{c})_{*}$
+for large enough $c$, equation 5.9 holds.
 Now we are ready to prove our main theorems. The main ingredient for each result is Equation 5.9.
 Theorem 5.11.
-Let $K\subset Y$ be a blow-up unknotted knot, and suppose $0<r<s$ are distinct positive rational numbers with the property that $Y_{r}(K)\cong\pm Y_{s}(K)$. Then $r,s>1$ or $K$ is the unknot.
+Let $K\subset Y$ be a blow-up bounding knot, and suppose $0<r<s\leq 1$ are distinct positive rational numbers with the property that $Y_{r}(K)\cong\pm Y_{s}(K)$.
+Then $K$ is the unknot.
+Furthermore, if $K\subset Y$ is a blow-up unknotted knot and $0<r<s$ are distinct positive rational numbers such that $Y_{r}(K)\cong\pm Y_{s}(K)$, then $r,s>1$ or $K$ is the unknot.
 Proof.
+Let $K\subset Y$ be blow-up bounding.
 Since $H_{1}(Y_{\pm p/q}(K))\cong H_{1}(Y)\oplus{\mathbb{Z}}/p{\mathbb{Z}},$ we can fix $p>0$ throughout.
 First, let $q$ be relatively prime to $p$ and $q\geq p$.
-Then $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}=\operatorname{Im}(\widehat{v}_{0})_{*}$ for all $j=0,\dots,p-1$.
-Thus $*t_{K}^{p/q}=p\operatorname{rank}(\widehat{v}_{0})_{*}$.
+Then $\operatorname{Im}(\widehat{v}_{j/q})_{*}\cap\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}=\operatorname{Im}(\widehat{v}_{0})_{*}$ for all $j=0,\dots,p-1$.
+Thus $t_{K}^{p/q}=p\operatorname{rank}(\widehat{v}_{0})_{*}$.
 Therefore, according to equation 5.9, for positive integral $q$ with $0<p/q\leq 1$, $\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ is a nondecreasing function of $q$.
@@
 Hence, by Proposition 5.5, $K$ is the unknot.
-Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result is true whenever $0<r<s\leq 1$.
-Now suppose $q,q^{\prime}$ are relatively prime to $p$ and $0<q<p<q^{\prime}$. Then
-$t_{K}^{p/q}-t_{K}^{p/q^{\prime}}$ $*\displaystyle\leq\sum_{j=0}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-p\operatorname{rank}(\widehat{v}_{0})_{*}$
-$*\displaystyle=\sum_{j=q}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-(p-q)\operatorname{rank}(\widehat{v}_{0})_{*}$
-$\leq(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*}).$
-Therefore,
+Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the first part of the theorem is true.
+Also, the second part is true for $0<r<s\leq 1$ since blow-up unknotted knots are blow-up bounding.
+Now suppose $K$ is blow-up unknotted and $q,q^{\prime}$ are relatively prime to $p$ and $0<q<p<q^{\prime}$. Then, with $)b=\operatorname{rank}(H_{*}(\widehat{B}$)),
+$\begin{split}t_{K}^{p/q}-t_{K}^{p/q^{\prime}}&\leq\sum_{j=0}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-p\operatorname{rank}(\widehat{v}_{0})_{*}\\ &=\sum_{j=q}^{p-1}\operatorname{rank}(\widehat{v}_{j/q})_{*}-(p-q)\operatorname{rank}(\widehat{v}_{0})_{*}\\ &\leq(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})\\ &=0\end{split}$ (5.12)
+since $(\widehat{v}_{0})_{*}$ is surjective by Lemma 5.4. By equation 5.9,
 $\operatorname{rank}\widehat{HF}(Y_{p/q^{\prime}}(K))$ $-\operatorname{rank}\widehat{HF}(Y_{p/q}(K))=(q^{\prime}-q)(\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*})$
 $+(q^{\prime}-q)\left(2\sum_{s=1}^{g-1}(\operatorname{rank}\ker(\widehat{v}_{s})_{*}+b-\operatorname{rank}(\widehat{v}_{s})_{*})\right)$
-$+2t_{K}^{p/q^{\prime}}-2t_{K}^{p/q}$
-$\geq 2(q^{\prime}-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})-2(p-q)(b-\operatorname{rank}(\widehat{v}_{0})_{*})$
+$+2t_{K}^{p/q^{\prime}}-2t_{K}^{p/q}.$
+The entire sum from $s=1$ to g1g-1 on the right hand side is nonnegative, so we ignore this term, and by Lemma 5.4, we have
+$\operatorname{rank}\ker(\widehat{v}_{0})_{*}+b-\operatorname{rank}(\widehat{v}_{0})_{*}=\operatorname{rank}\ker(\widehat{v}_{0})_{*}.$
+Together with inequality 5.12, we get
+$\operatorname{rank}\widehat{HF}(Y_{p/q^{\prime}}(K))-\operatorname{rank}\widehat{HF}(Y_{p/q}(K))$ $\geq(q^{\prime}-q)\operatorname{rank}\ker(\widehat{v}_{0})_{*}$
 $\geq 0,$
@@
 Theorem 5.14.
-Let $K\subset Y$ be a nontrivial blow-up unknotted knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$.
-Furthermore, if $g(K)\geq 2$ or $\operatorname{rank}H_{*}(\widehat{A}_{0})>b$, then $Y_{1}(K)\ncong\pm Y$.
+Let $K\subset Y$ be a nontrivial blow-up bounding knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $K$ is a nontrivial blow-up unknotted knot, then $Y_{1}(K)\ncong\pm Y$.
 Proof.
@@
 If $q=1$, then
-$\operatorname{rank}\widehat{HF}(Y_{1}(K))\geq\operatorname{rank}(\widehat{v}_{0})_{*}+\operatorname{rank}\ker(\widehat{v}_{0})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{0})\geq b,$
-and it is strictly greater unless $\operatorname{rank}H_{*}(\widehat{A}_{0})=b$ and $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 1$, which implies $g(K)\leq 1$.
+$\operatorname{rank}\widehat{HF}(Y_{1}(K))\geq\operatorname{rank}(\widehat{v}_{0})_{*}+\operatorname{rank}\ker(\widehat{v}_{0})_{*}=\operatorname{rank}H_{*}(\widehat{A}_{0})\geq b$
+by Lemma 5.4 since $K$ is blow-up unknotted, and it is strictly greater unless $(\widehat{v}_{s})_{*}$ is an isomorphism for all $s\geq 0$, which implies $K$ is the unknot.
 Since the total rank of $\widehat{HF}$ is an invariant of the underlying (unoriented) three-manifold, the result holds. ∎
+Acknowledgements The author would like to thank Yi Ni for advising him in this project and the referee for many helpful comments.